### v1

# Parameterized Controls and Isomorphic Representations of G-Expectations

Abstract
This paper establishes a fundamental isomorphism between the abstract framework of $G$-expectation and a concrete parameterized family of classical stochastic processes, providing a unified framework for nonlinear stochastic calculus under model uncertainty. We demonstrate that $G$-Brownian motion and related stochastic differential equations (SDEs) can be intrinsically represented as controlled processes $\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$, where $\theta$ denotes an admissible control process taking values in a compact convex parameter set $\Theta$ encoding volatility uncertainty.
The core contribution is the construction of a complete algebraic and analytic isomorphism $\Phi$ that preserves all operations including addition, multiplication, function composition, and limit transitions. This isomorphism provides a rigorous mathematical foundation for understanding $G$-martingales, with a distinction between symmetric martingales (fair in all volatility scenarios) and asymmetric martingales (favorable but fair only in the worst-case scenario).
We develop a comprehensive theory for solving $G$-SDEs through their parameterized representations, establishing existence and uniqueness results under standard Lipschitz and growth conditions. The framework reveals how bounded control sets preserve these classical conditions while capturing volatility uncertainty.
The parameterized control perspective offers both mathematical rigor and computational practicality, enabling the transfer of classical stochastic calculus results to the $G$-expectation framework while maintaining the essential sublinear structure of model uncertainty.
Key Words: G-Expectation, G-Brownian Motion, Parameterized Controls, Isomorphic Representations, Volatility Uncertainty, Dynamic Programming Principle
E-mail: zhaoguangqian@mail.ustc.edu.cn
1 Introduction
The quantification and management of model uncertainty—often referred to as Knightian uncertainty or ambiguity—represents a fundamental challenge across several disciplines, most notably in mathematical finance and decision theory. Traditional stochastic calculus, built upon the foundation of a single, precisely known probability measure, is inherently limited in its capacity to address situations where the model itself is imperfectly known. This limitation is particularly acute in financial markets, where volatility—the key parameter in option pricing and risk management—is notoriously difficult to estimate and is subject to significant time-varying uncertainty.
The groundbreaking framework of $G$-expectation, pioneered by Peng , provides a rigorous mathematical foundation for modeling such uncertainty. By replacing the classical linear expectation with a sublinear operator, the $G$-framework successfully captures the essence of volatility ambiguity. Within this framework, the $G$-Brownian motion $B_{t}$ and related stochastic differential equations ($G$-SDEs) have been defined, with their quadratic variation $\langle B\rangle_{t}$ evolving within a prescribed interval $\underline{\sigma}^{2}t,\overline{\sigma}^{2}t]$, naturally encoding the uncertainty in volatility. This leads to a rich theory of nonlinear stochastic calculus .
A profound insight emerging from this framework is the intrinsic connection between $G$-expectations and stochastic control theory. Specifically, the $G$-expectation can be naturally represented as the value function of a stochastic control problem:
$\hat{\mathbb{E}}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right],$
where the control process $\theta_{s}$ takes values in a compact set $\Theta$ representing the volatility uncertainty. This representation naturally leads to a Hamilton-Jacobi-Bellman (HJB) equation through the Dynamic Programming Principle (DPP):
$\partial_{t}v+G(D_{x}^{2}v)=0,\quad v(T,x)=\varphi(x),$
where the nonlinearity $G(A)=\frac{1}{2}\sup_{\gamma\in\Theta}\text{tr}(A\gamma\gamma^{\top})$ encodes the uncertainty set. The DPP manifests infinitesimally as:
$v(t,x)=\sup_{\theta\in\mathcal{A}_{t,t+h}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[v\left(t+h,x+\int_{t}^{t+h}\theta_{s}dW_{s}\right)\right].$
Despite this elegant connection, the abstract nature of the $G$-expectation can present a barrier to intuitive understanding and practical application. A central question thus arises:
Can we establish a more explicit and operational connection between the abstract $G$-expectation space and the concrete framework of parameterized controlled processes?
This paper provides a definitive affirmative answer by constructing a complete isomorphism between these two domains. We demonstrate that the entire $G$-expectation framework admits a rigorous correspondence to a family of classical stochastic processes controlled by volatility parameters. The core idea is to represent a $G$-Brownian motion path as a family of classical Itô processes:
$\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}},\quad\text{where}\quad B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}.$
The sublinear $G$-expectation is then revealed to be the supremum over all these classical scenarios, providing a stochastic control interpretation that complements the PDE perspective.
This isomorphic correspondence, denoted by $\Phi$, is shown to preserve the entire algebraic and analytic structure—including addition, multiplication, function composition, and limit transitions—thereby bridging the abstract nonlinear theory with the well-understood classical theory.
Contributions and Outline: The main contributions of this paper are as follows:
  • We construct a precise isomorphism $\Phi$ between the $G$-expectation space and a space of parameterized controlled processes, providing a concrete representation that complements the HJB equation characterization (Section 3).
  • We develop a comprehensive theory for $G$-SDEs, proving existence and uniqueness of solutions through their parameterized representations, thereby connecting the SDE solutions with the dynamic programming framework (Section 4).
  • Through this isomorphism, we characterize the fundamental structure of $G$-martingales, introducing a distinction between symmetric martingales (fair in all volatility scenarios) and asymmetric martingales (favorable but fair only in the worst-case scenario) (Sections 5 .
  • We resolve subtle definability difficulties associated with conditional $G$-expectations, clarifying the interplay between supremum operations and conditional expectations under model uncertainty.
The paper is organized as follows: Section 2 lays the theoretical foundation, formalizing $G$-expectation as a value function and deriving its connection to dynamic programming and nonlinear PDEs. Section 3 introduces the fundamental isomorphism, the central concept of this work. Sections 4 and 5 delve into the solutions of $G$-SDEs and the properties of $G$-martingales from this new perspective, respectively. Finally, Section 6 concludes.
Our work not only deepens the theoretical understanding of the $G$-framework but also unlocks new computational pathways. By translating complex nonlinear problems into families of more tractable classical ones, this parameterized control perspective offers a powerful tool for both theoretical analysis and practical computation in the presence of model uncertainty, while maintaining a natural connection to the HJB and dynamic programming framework.
2 Theoretical Foundation
2.1 G-Expectation as a Value Function
The fundamental insight is that G-expectation can be represented as the value function of a stochastic control problem. For a G-Brownian motion $B_{t}$ and terminal payoff $\varphi\in C_{b}(\mathbb{R}^{d})$, we have:
$\hat{\mathbb{E}}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}^{\Theta}_{0,T}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right]$ (2.1)
where:
  • $\mathcal{A}^{\Theta}_{0,T}$ denotes admissible controls taking values in $\Theta$
  • $W_{t}$ is a classical Brownian motion under reference measure $\mathbb{P}$
  • $\Theta=[\underline{\sigma}^{2}I_{d},\overline{\sigma}^{2}I_{d}]\subset S^{d}_{+}$ models volatility uncertainty
This equivalence reveals that:
  1. 1.
    Volatility ambiguity is characterized by the control set $\Theta$
  2. 2.
    G-Brownian paths emerge as controlled processes: $B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}$
2.2 Dynamic Programming and Nonlinear PDEs
The control-theoretic perspective leads to a Hamilton-Jacobi-Bellman (HJB) equation. Define the value function:
$v(t,x)=\sup_{\theta\in\mathcal{A}^{\Theta}_{t,T}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(x+\int_{t}^{T}\theta_{s}dW_{s}\right)\right]$ (2.2)
It satisfies the nonlinear parabolic PDE:
$\begin{cases}\partial_{t}v+G(D_{x}^{2}v)=0,&(t,x)\in[0,T)\times\mathbb{R}^{d}\\ v(T,x)=\varphi(x)\end{cases}$ (2.3)
where the nonlinearity $G$ encodes the uncertainty set:
$G(A)=\frac{1}{2}\sup_{\gamma\in\Theta}\text{tr}(A\gamma\gamma^{\top}),\quad A\in S^{d}$ (2.4)
This PDE is the infinitesimal manifestation of the Dynamic Programming Principle (DPP):
$v(t,x)=\sup_{\theta\in\mathcal{A}^{\Theta}_{t,t+h}}\mathbb{E}_{\mathbb{P}}\left[v\left(t+h,x+\int_{t}^{t+h}\theta_{s}dW_{s}\right)\right]$ (2.5)
3 The Fundamental Isomorphism: G-Brownian motion as Parameterized Controlled Processes
The central insight of this work is establishing a natural isomorphism between elements in the G-expectation space and parameterized families of controlled processes. Specifically:
  1. 1.
    G-Brownian motion as controlled paths: For each volatility control process $\theta\in\Theta$, we define:
    $B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}$
    with quadratic variation correctly given by:
    $d\langle B^{\theta}\rangle_{t}^{ij}=(\theta_{t}\theta_{t}^{\top})^{ij}dt$
  2. 2.
    G-SDE solutions as parameterized families: For a general G-SDE:
    $dX_{t}=b(X_{t})dt+h_{ij}(X_{t})d\langle B\rangle_{t}^{ij}+\sigma_{j}(X_{t})dB_{t}^{j}$
    the solution is intrinsically represented by the parameterized family:
    $X_{t}^{\theta}=X_{0}+\int_{0}^{t}b(X_{s}^{\theta})ds+\int_{0}^{t}h_{ij}(X_{s}^{\theta})(\theta_{s}\theta_{s}^{\top})^{ij}ds+\int_{0}^{t}\sigma_{j}(X_{s}^{\theta})\theta_{s}^{j}dW_{s}^{j}$
  3. 3.
    The isomorphism: There exists a bijective correspondence:
    $\begin{array}[]{c}\text{G-expectation space}\\ \Updownarrow\\ \left\{X^{\theta}:\theta\in\Theta\right\}\end{array}$
    where the parameterized family $\{X^{\theta}\}_{\theta\in\Theta}$ is the G-SDE solution
This isomorphism reveals that:
  • G-processes are intrinsically controlled trajectories parameterized by volatility uncertainty
  • The G-expectation framework provides a geometric encoding of model uncertainty
  • Solutions to G-SDEs emerge as universal families of classical SDE solutions
3.1 Classical G-Expectation and G-Brownian Motion
3.1.1 G-Normal Distribution
The cornerstone of the theory is the G-normal distribution . Let $\xi$ be a random variable. It is said to follow a G-normal distribution, denoted $\xi\sim\mathcal{N}(0,[\underline{\sigma}^{2},\overline{\sigma}^{2}])$, if for any $\phi\in C_{b}(\mathbb{R})$, its sublinear expectation is defined by the following nonlinear heat equation:
$\mathbb{E}[\phi(\xi)]=u(1,0),$ (3.1)
where $u(t,x)$ is the unique viscosity solution to the following nonlinear partial differential equation (PDE):
$\frac{\partial u}{\partial t}$ $=G\left(\frac{\partial^{2}u}{\partial x^{2}}\right),\quad(t,x)\in(0,\infty)\times\mathbb{R},$ (3.2)
$u(0,x)$ $=\phi(x),$ (3.3)
with the nonlinear generator $G:\mathbb{R}\to\mathbb{R}$ given by:
$G(\alpha)=\frac{1}{2}\sup_{\sigma\in[\underline{\sigma},\overline{\sigma}]}(\sigma^{2}\alpha)=\frac{1}{2}\left(\overline{\sigma}^{2}\alpha^{+}-\underline{\sigma}^{2}\alpha^{-}\right).$ (3.4)
Here, $\alpha^{+}=\max(\alpha,0)$ and $\alpha^{-}=-\min(\alpha,0)$. This distribution is normal in the sense that it is centered and stable under scaling and addition of independent random variables, yet it is not linear due to the sublinearity of $G$.
A particularly revealing special case arises when the uncertainty set collapses to a single point, i.e., $\underline{\sigma}=\overline{\sigma}=\sigma_{0}>0$. In this case, the generator becomes linear, $G(\alpha)=\frac{1}{2}\sigma_{0}^{2}\alpha$, and the G-normal distribution reduces to the classical normal distribution $\mathcal{N}(0,\sigma_{0}^{2})$.
3.1.2 G-Brownian Motion
[…]
  1. (i)
    $B_{0}=0$.
  2. (ii)
    For any $t,s\geq 0$, the increment $B_{t+s}-B_{t}$ is identically distributed to $B_{s}$, and $B_{t+s}-B_{t}$ is independent of $(B_{t_{1}},B_{t_{2}},\dots,B_{t_{n}})$ for any $n\in\mathbb{N}$ and $0\leq t_{1}\leq\dots\leq t_{n}\leq t$.
  3. (iii)
    $\lim_{t\downarrow 0}\hat{E}[|B_{t}|^{3}]t^{-1}=0$.
Moreover, if $\hat{E}[B_{t}]=\hat{E}[-B_{t}]=0$ for all $t$, then $(B_{t})_{t\geq 0}$ is called a symmetric G-Brownian motion.
In the context of G-expectation, the increments of G-Brownian motion follow a G-normal distribution. Specifically, for any $t,s\geq 0$, $B_{t+s}-B_{t}\sim\mathcal{N}(0,[\underline{\sigma}^{2}s,\overline{\sigma}^{2}s])$, where $\mathcal{N}(0,[\underline{\sigma}^{2},\overline{\sigma}^{2}])$ denotes the G-normal distribution with uncertainty interval $\underline{\sigma}^{2},\overline{\sigma}^{2}]$.
The quadratic variation process $\langle B\rangle_{t}$ of G-Brownian motion is a continuous, non-decreasing process such that ,
$\underline{\sigma}^{2}t\leq\langle B\rangle_{t}\leq\overline{\sigma}^{2}t.$
This process captures the accumulated uncertainty over time.
The corresponding infinitesimal generator for a function $\phi$ is given by the same nonlinear operator $G$:
$\lim_{t\downarrow 0}\frac{\mathbb{E}[\phi(x+B_{t})]-\phi(x)}{t}=G(\phi^{\prime\prime}(x)).$ (3.5)
3.2 A Generalized Construction via Random Variables
This generator governs the evolution of expectations under the G-Brownian motion. The drift term is zero, reflecting the martingale property of $B_{t}$, and the diffusion term is given by the supremum over constant parameters. This establishes the foundational link: classical G-Brownian motion corresponds to a constant diffusion coefficient uncertainty set $\Theta=[\underline{\sigma},\overline{\sigma}]$.
Definition 3.1 (Construction via Supremum over Linear Expectations).
Let $(\Omega,\mathcal{F},P)$ be a probability space sufficient to support a standard Brownian motion $\{W_{t}\}_{t\geq 0}$. For a given compact, convex parameter set $\Lambda\subset\mathbb{R}^{n}$, we define two fundamental random objects:
1. Max-Distribution: For a parameter $\mu\in\Lambda$, define the random variable $\eta(\mu)=\mu$. The associated nonlinear expectation $\mathbb{E}[\eta]$ is defined for a test function $\phi$ as:
$\mathbb{E}[\phi(\eta)]:=\sup_{\mu\in\Lambda}E_{P}[\phi(\eta(\mu))].$ (3.6)
This object captures the pure parameter uncertainty in its most extreme form.
2. G-Normal Distribution: For a parameter $\sigma\in\Lambda$, define the random variable $X(\sigma)=\sigma W_{1}$ (which is classically normal distributed, $X(\sigma)\sim N(0,\sigma^{2})$). The associated nonlinear expectation $\mathbb{E}[X]$ is defined as:
$\mathbb{E}[\phi(X)]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(X(\sigma))]=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{1})].$ (3.7)
This constitutes a G-normal distribution with the uncertainty set $\Lambda$.
3. Quasi G-Brownian Motion: We define a process $\{B_{t}\}_{t\geq 0}$ by its action on test functions. For any $\phi\in C_{l,Lip}(\mathbb{R})$, its nonlinear expectation is given by:
$\mathbb{E}[\phi(B_{t})]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t})].$ (3.8)
Furthermore, for a process of the form $H_{t}=bB_{t}+mt$ (where the drift $m$ is also subject to uncertainty within a convex set), the expectation is defined as:
$\mathbb{E}[\phi(H_{t})]:=\sup_{(b,m)\in\Lambda^{\prime}}E_{P}[\phi(bW_{t}+mt)],$ (3.9)
where $\Lambda^{\prime}$ is a compact convex set in $\mathbb{R}^{2}$.
These constructions are defined on a filtered probability space represented by the quadruple:
$(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{E})$ (3.10)
where:
  • $\Omega$ is the sample space, representing the set of all possible outcomes;
  • $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$, representing the set of all measurable events;
  • $(\mathcal{F}_{t})_{t\geq 0}$ is a filtration, i.e., an increasing family of sub-$\sigma$-algebras of $\mathcal{F}$ satisfying $\mathcal{F}_{s}\subseteq\mathcal{F}_{t}$ for all $0\leq s\leq t$, which represents the accumulation of information over time;
  • $\mathbb{E}$ is the sublinear expectation operator defined in equations (3.6) and (3.7), which generalizes the classical linear expectation to account for model uncertainty.
Remark 3.1.
The construction of the ”Quasi G-Brownian Motion” and the related objects herein represents the core methodological approach of this paper. It is intended to schematically illustrate the central idea of handling model uncertainty via the supremum over linear expectations. It is important to note that the process $\{B_{t}\}$ defined here is for methodological demonstration purposes; while it shares the same one-dimensional distributions with the strict mathematical G-Brownian motion, their joint distributions are entirely different, and thus they are not mathematically equivalent.
Definition 3.2 (General Construction of Sublinear Expectation via Model Uncertainty).
Let $(\Omega,\mathcal{F},P)$ be a classical probability space. Let $\Theta$ be a non-empty, compact, and convex parameter set representing the domain of model uncertainty.
Consider a generalparameter-controlled random variable $X(\theta)$, where $\theta\in\Theta$ is the parameter that dictates the model specification. The associated sublinear expectation $\mathbb{E}[\cdot]$ is defined for a test function $\phi$ (typically in the space of Lipschitz continuous functions $C_{l,Lip}(\mathbb{R}^{d})$) by the following operation:
$\mathbb{E}[\phi(X)]:=\sup_{\theta\in\Theta}E_{P}[\phi(X(\theta))].$ (3.11)
This means the operator $\mathbb{E}$ acts on the random variable $X$ by a two-step procedure:
1. For a fixed parameter $\theta\in\Theta$, compute the classical linear expectation $E_{P}$ of the random variable $\phi(X(\theta))$ under the probability measure $P$.
2. Subsequently, take the supremum of these resulting real numbers over all admissible parameters $\theta$ in the uncertainty set $\Theta$.
This operator is defined on the sublinear expectation space:
$(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{E})$ (3.12)
where:
  • $\Omega$ is the sample space;
  • $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$;
  • $(\mathcal{F}_{t})_{t\geq 0}$ is a filtration representing the information flow over time;
  • $\mathbb{E}$ is the sublinear expectation operator defined above.
Remarks:
  • This general definition subsumes the specific constructions (e.g., Max-Distribution, G-Normal Distribution, Generalized G-Brownian Motion) presented in Definition 3.1 as special cases by choosing the parameter-controlled variable $X(\theta)$ and the uncertainty set $\Theta$ appropriately.
  • The operator $\mathbb{E}$ defined in this way is sublinear: $\mathbb{E}[X+Y]\leq\mathbb{E}[X]+\mathbb{E}[Y]$ and $\mathbb{E}[\lambda X]=\lambda\mathbb{E}[X]$ for $\lambda\geq 0$. It also preserves constants: $\mathbb{E}[c]=c$ for any constant $c$.
While the parameterized construction provides valuable insight, it is crucial to recognize its fundamental limitation: constant control sets cannot capture the path-dependent nature of true G-Brownian motion.
Consider the parameterized ”quasi G-Brownian motion” defined by:
$B_{t}^{\sigma}=\sigma W_{t},\quad\sigma\in\Lambda$ (3.13)
with nonlinear expectation:
$\mathbb{E}[\phi(B_{t})]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t})]$ (3.14)
Key observations:
  1. 1.
    One-dimensional distributions match: For fixed $t$,
    $\text{Law}_{G}(B_{t})\cong\sup_{\sigma\in\Lambda}\text{Law}_{P}(\sigma W_{t})$
    meaning marginal distributions coincide with true G-Brownian motion.
  2. 2.
    Path structure diverges: The joint distributions differ fundamentally:
    $\mathbb{E}[\phi(B_{t},B_{s})]=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t},\sigma W_{s})]\neq\hat{\mathbb{E}}[\phi(B_{t},B_{s})]$
    This occurs because constant $\sigma$ cannot capture the time-consistent volatility uncertainty of true G-Brownian motion.
  3. 3.
    The Difficulty in Defining Conditional Expectation
3.3 True G-Brownian
To overcome these limitations, we must introduce time-varying adaptive controls:
Definition 3.3 (True G-Brownian Motion via Adaptive Controls).
Let $\Theta\subset\mathbb{R}^{d\times d}$ be compact convex. Define:
$B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s},\quad\theta\in\mathcal{A}^{\Theta}$ (3.15)
with quadratic variation:
$d\langle B^{\theta}\rangle_{t}=(\theta_{t}\theta_{t}^{\top})dt$ (3.16)
The genuine G-Brownian motion is then characterized by:
${\mathbb{E}}[\phi(B_{t})]=\sup_{\theta\in\mathcal{A}^{\Theta}}E_{P}[\phi(B_{t}^{\theta})]$ (3.17)
The space of admissible control processes is defined as:
$\mathcal{A}^{\Theta}=\left\{\theta:[0,\infty)\times\Omega\to\Theta\ \middle|\ \theta\text{ is }\mathcal{F}_{t}\text{-adapted and }\mathbb{E}\left[\int_{0}^{T}|\theta_{t}|^{2}dt\right]<\infty\ \forall T>0\right\}$
where $\Theta\subset\mathbb{R}^{d\times d}$ is a compact convex set representing the volatility uncertainty.
For the G-Brownian motion $B_{t}$, the conditional expectation is:
${\mathbb{E}}_{t}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}_{t,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{t}^{T}\theta_{s}dW_{s}\right)\middle|\mathcal{F}_{t}\right]$
This definition provides the mathematical foundation for conditional expectations in the G-framework, capturing the essence of model uncertainty through the supremum over volatility controls.
Definition 3.4 (Conditional G-Expectation Operator).
For a G-Brownian motion $\{B_{t}\}_{t\geq 0}$ and a test function $\varphi\in C_{b.Lip}(\mathbb{R}^{d})$, the conditional G-expectation operator $\hat{\mathbb{E}}_{s}[\cdot]$ is defined as:
${\mathbb{E}}_{s}[\varphi(B_{t})]=\sup_{\theta\in\mathcal{A}_{s,t}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{s}+\int_{s}^{t}\theta_{u}dW_{u}\right)\middle|\mathcal{F}_{s}\right]$
where:
  • $\mathcal{A}_{s,t}^{\Theta}$ denotes admissible control processes on $[s,t]$
  • $\theta_{u}\in\Theta=[\underline{\sigma},\overline{\sigma}]$ represents volatility controls
  • $W_{u}$ is a standard Brownian motion under $\mathbb{P}$
Definition 3.5 (Conditional G-Expectation for Cylinder Functions).
For any cylinder function of the form:
$\xi=\varphi(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}}),\quad\varphi\in C_{b.Lip}(\mathbb{R}^{d\times n})$
with $0\leq t_{1}<t_{2}<\cdots<t_{n}\leq T$, the conditional G-expectation is defined as:
${\mathbb{E}}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\middle|\mathcal{F}_{s}\right]$
where $B_{t}^{\theta}=\int_{0}^{t}\theta_{u}dW_{u}$ is the controlled process.
This definition coincides with the classical G-expectation definition when $s=0$:
${\mathbb{E}}[\xi]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right]$
Theorem 3.6 (Conditional Expectation Simplification).
For any cylinder function $\xi=\varphi(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}})$ and $s\leq T$, the conditional G-expectation simplifies to:
${\mathbb{E}}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(x_{1},x_{2},\cdots,x_{k-1},B_{t_{k}}^{\theta}-B_{t_{k-1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right]\bigg{|}_{x_{i}=B_{t_{i}}}$
where for $t_{k-1}\leq s$$t_{k}>s$, we treat $B_{t_{i}}$ as constants $x_{i}$ in the expectation.
Note that our definition of conditional expectation here is completely consistent with the classical definition of G-conditional expectation under the definition of cylindrical Brownian motion, so we can similarly extend it to its completion space.
Theorem 3.7 (Fundamental Properties).
The conditional G-expectation operator satisfies:
  1. 1.
    Sublinearity: ${\mathbb{E}}_{s}[X+Y]\leq{\mathbb{E}}_{s}[X]+{\mathbb{E}}_{s}[Y]$
  2. 2.
    Positive Homogeneity: ${\mathbb{E}}_{s}[\lambda X]=\lambda{\mathbb{E}}_{s}[X]$ for $\lambda\geq 0$
  3. 3.
    Preservation of Constants: ${\mathbb{E}}_{s}[c]=c$ for any constant $c$
  4. 4.
    Monotonicity: If $X\leq Y$ , then ${\mathbb{E}}_{s}[X]\leq{\mathbb{E}}_{s}[Y]$
Theorem 3.8 (Tower Property).
For $s\leq t\leq u$, the conditional expectation satisfies:
${\mathbb{E}}_{s}\left[{\mathbb{E}}_{t}[X]\right]={\mathbb{E}}_{s}[X]$
This ensures the time consistency of the G-expectation framework.
3.4 The Complete Isomorphism
To establish a full isomorphism, we must introduce time-varying adaptive controls.
3.4.1 Adaptive Control Representation
Definition 3.9 (True G-Brownian Motion via Adaptive Controls).
Let $\Theta\subset\mathbb{R}$ be compact convex. Define:
$B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s},\quad\theta\in\mathcal{A}^{\Theta}$ (3.18)
with quadratic variation:
$d\langle B^{\theta}\rangle_{t}=\theta_{t}^{2}dt$ (3.19)
The genuine G-Brownian motion is characterized by:
$\hat{\mathbb{E}}[\phi(B_{t})]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[\phi(B_{t}^{\theta})]$ (3.20)
3.4.2 Algebraic Structure Preservation
The isomorphism preserves all algebraic operations:
Theorem 3.10 (Algebraic Isomorphism).
The mapping $\Phi:X\mapsto\{X^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ preserves:
  1. 1.
    Addition: $\Phi(X+Y)=\Phi(X)+\Phi(Y)$
  2. 2.
    Scalar Multiplication: $\Phi(\lambda X)=\lambda\Phi(X)$
  3. 3.
    Multiplication: $\Phi(X\cdot Y)=\Phi(X)\cdot\Phi(Y)$
  4. 4.
    Function Composition: $\Phi(\phi(X))=\phi(\Phi(X))$
The preservation follows directly from the pathwise definition:
$(X+Y)^{\theta}$ $=X^{\theta}+Y^{\theta}$
$(\lambda X)^{\theta}$ $=\lambda X^{\theta}$
$(X\cdot Y)^{\theta}$ $=X^{\theta}\cdot Y^{\theta}$
$\phi(X)^{\theta}$ $=\phi(X^{\theta})$
Fourth, we define the multivariate composition in a similar manner. Building upon the correspondence established via G-Brownian motion, we derive a corresponding relationship in the space spanned by these two elements, while ensuring that the mapping constitutes a bijection.
3.4.3 Analytic Structure Preservation
The isomorphism also preserves analytic properties:
Theorem 3.11 (Analytic Isomorphism).
The mapping $\Phi$ preserves:
  1. 1.
    Expectation: $\hat{\mathbb{E}}[X]=\sup_{\theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
  2. 2.
    Conditional Expectation: $\hat{\mathbb{E}}_{t}[X]=\sup_{\theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}|\mathcal{F}_{t}]$
  3. 3.
    Convergence: $X_{n}\to X$ in $L_{G}^{p}$ iff $X_{n}^{\theta}\to X^{\theta}$ in $L^{p}$ uniformly over $\theta$
3.5 Stochastic Calculus Equivalence
The isomorphism extends to differential operations:
3.5.1 Itô Calculus Preservation
Theorem 3.12 (Itô Calculus Isomorphism).
For G-Itô processes:
$dX_{t}$ $=b_{t}dt+h_{t}d\langle B\rangle_{t}+\sigma_{t}dB_{t}$
$dX_{t}^{\theta}$ $=b_{t}^{\theta}dt+h_{t}^{\theta}\theta_{t}^{2}dt+\sigma_{t}^{\theta}\theta_{t}dW_{t}$
we have:
  1. 1.
    $\Phi(dB_{t})=\{\theta_{t}dW_{t}\}$
  2. 2.
    $\Phi(\langle B\rangle_{t})=\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}$
  3. 3.
    $\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}$
3.5.2 Itô Calculus Isomorphism (Integral Form)
Proof.
We prove each component using integral representations and L²-convergence:
Part 1: $\Phi\left(\int_{0}^{t}dB_{s}\right)=\left\{\int_{0}^{t}\theta_{s}dW_{s}\right\}$
By definition of the isomorphism:
$\Phi\left(\int_{0}^{t}dB_{s}\right)=\Phi(B_{t})=\left\{B_{t}^{\theta}\right\}_{\theta\in\mathcal{A}^{\Theta}}=\left\{\int_{0}^{t}\theta_{s}dW_{s}\right\}_{\theta\in\mathcal{A}^{\Theta}}$
This holds identically as:
$B_{t}=\int_{0}^{t}dB_{s},\quad B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}$
Part 2: $\Phi\left(\langle B\rangle_{t}\right)=\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}$
The quadratic variation is defined as the L²-limit:
$\langle B\rangle_{t}=\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}(B_{t_{k}}-B_{t_{k-1}})^{2}\quad\text{in }L_{G}^{2}$
Applying $\Phi$:
$\Phi(\langle B\rangle_{t})=\Phi\left(\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}(B_{t_{k}}-B_{t_{k-1}})^{2}\right)$
Since $\Phi$ preserves L²-convergence:
$\Phi(\langle B\rangle_{t})=\left\{\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}(B_{t_{k}}^{\theta}-B_{t_{k-1}}^{\theta})^{2}\right\}_{\theta\in\mathcal{A}^{\Theta}}\quad\text{in }L^{2}$
For each fixed $\theta$, by Itô isometry:
$\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}\left(\int_{t_{k-1}}^{t_{k}}\theta_{s}dW_{s}\right)^{2}=\int_{0}^{t}\theta_{s}^{2}ds\quad\text{in }L^{2}(\Omega,\mathbb{P})$
Thus:
$\Phi(\langle B\rangle_{t})=\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}_{\theta\in\mathcal{A}^{\Theta}}$
Part 3: $\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}$
The G-Itô integral is defined as:
$\int_{0}^{t}H_{s}dB_{s}=\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}H_{t_{k-1}}(B_{t_{k}}-B_{t_{k-1}})\quad\text{in }L_{G}^{2}$
Applying $\Phi$:
$\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}H_{t_{k-1}}^{\theta}(B_{t_{k}}^{\theta}-B_{t_{k-1}}^{\theta})\right\}_{\theta\in\mathcal{A}^{\Theta}}$
For each fixed $\theta$, the classical Itô integral satisfies:
$\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}H_{t_{k-1}}^{\theta}\int_{t_{k-1}}^{t_{k}}\theta_{s}dW_{s}=\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\quad\text{in }L^{2}(\Omega,\mathbb{P})$
Thus:
$\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}_{\theta\in\mathcal{A}^{\Theta}}$
Consistency with G-Itô Formula
Consider the integral form of G-Itô formula for $f(B_{t})$:
$f(B_{t})-f(0)=\int_{0}^{t}\partial_{x}f(B_{s})dB_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s})d\langle B\rangle_{s}$
Applying $\Phi$:
$\Phi(f(B_{t})-f(0))=\left\{f(B_{t}^{\theta})-f(0)\right\}_{\theta\in\mathcal{A}^{\Theta}}$
By classical Itô formula for each $\theta$:
$f(B_{t}^{\theta})-f(0)=\int_{0}^{t}\partial_{x}f(B_{s}^{\theta})\theta_{s}dW_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s}^{\theta})\theta_{s}^{2}ds$
Thus:
$\Phi\left(\int_{0}^{t}\partial_{x}f(B_{s})dB_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s})d\langle B\rangle_{s}\right)=\left\{\int_{0}^{t}\partial_{x}f(B_{s}^{\theta})\theta_{s}dW_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s}^{\theta})\theta_{s}^{2}ds\right\}_{\theta\in\mathcal{A}^{\Theta}}$
This confirms the isomorphism preserves the integral structure of Itô calculus in L²-sense. ∎
3.6 Visualizing the Isomorphism: From G-SDE to Parameterized Processes
[…]
The isomorphism $\Phi$ establishes a complete correspondence between the G-expectation space and the parameterized classical space:
with the mapping defined as:
$\Phi(X)=\{X^{\theta}\}_{\theta\in\Theta}$
3.7 Algebraic and Analytic Structure Preservation
The isomorphism preserves all algebraic and analytic operations:
Operation G-Expectation Space Parameterized Space
Addition $X+Y$ $\Phi(X)+\Phi(Y)=\{X^{\theta}+Y^{\theta}\}$
Multiplication $X\cdot Y$ $\Phi(X)\cdot\Phi(Y)=\{X^{\theta}\cdot Y^{\theta}\}$
Scalar Multiplication $\lambda X$ $\lambda\Phi(X)=\{\lambda X^{\theta}\}$
Function Composition $\phi(X)$ $\phi(\Phi(X))=\{\phi(X^{\theta})\}$
Limit Operation $\lim\limits_{n\to\infty}X_{n}$ $\lim\limits_{n\to\infty}\Phi(X_{n})=\{\lim\limits_{n\to\infty}X_{n}^{\theta}\}$
Stochastic Integral $\int_{0}^{t}H_{s}dB_{s}$ $\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}$
Quadratic Variation $\langle B\rangle_{t}$ $\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}$
The sublinear expectation operators are related through the isomorphism:
G-Expectation Space Parameterized Space
$\hat{\mathbb{E}}[X]$ $\sup\limits_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
$\hat{\mathbb{E}}_{t}[X]$ $\sup\limits_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}\mid\mathcal{F}_{t}]$
  1. 1.
    Intrinsic Nature of Expectations:
    • $\hat{\mathbb{E}}$ is the intrinsic sublinear expectation operator of the G-expectation space
    • $\mathbb{E}_{\mathbb{P}}$ is the intrinsic linear expectation operator of the classical probability space
    • The supremum operation $\sup_{\theta}$ is the bridge connecting these intrinsic structures
  2. 2.
    sublinear expectation operators:
    $\hat{\mathbb{E}}[X]=\sup_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
    This formula does not define $\hat{\mathbb{E}}$, but rather expresses how the intrinsic G-expectation can be computed using the parameterized classical expectations.
  3. 3.
    Analytic Continuity: The isomorphism preserves convergence properties:
    $X_{n}\to X\text{ in }{L^{p}_{\mathbb{G}}}\iff X_{n}^{\theta}\to X^{\theta}\text{ in }L^{p}\text{ uniformly over }\theta$
This isomorphism provides:
  • A computational bridge from nonlinear to linear stochastic calculus
  • A theoretical foundation for model uncertainty quantification
  • An operational framework for solving G-SDEs through parameterized classical SDEs
4 How to Solve G-Stochastic Differential Equations
Consider a classical SDE parameterized by $\theta\in\Theta$:
$dX_{t}^{\theta}=b(t,X_{t}^{\theta})dt+\sigma(t,X_{t}^{\theta})\theta_{t}dW_{t}$ (4.1)
where:
  • $W_{t}$ is a standard Brownian motion
  • $\theta_{t}\in\Theta$ is an adapted control process
  • $\Theta\subset\mathbb{R}$ is compact and convex (bounded controls)
Under standard Lipschitz and growth conditions on $b$ and $\sigma$, for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$.
We reformulate the classical SDE as a G-SDE driven by G-Brownian motion $B_{t}$:
The classical SDE (4.1) is equivalent to the following G-SDE:
$dX_{t}=b(t,X_{t})dt+\sigma(t,X_{t})dB_{t}$ (4.2)
where $B_{t}$ is a G-Brownian motion
Theorem 4.1 (Existence and Uniqueness for G-SDE).
Under bounded controls $\theta_{t}\in\Theta$ and standard Lipschitz conditions:
$|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|$ $\leq K|x-y|$
$|b(t,x)|^{2}+|\sigma(t,x)|^{2}$ $\leq K(1+|x|^{2})$
the G-SDE (4.2) admits a unique solution in the quasi-sure sense:
$X_{t}=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
Proof.
The proof follows from:
  1. 1.
    For each fixed $\theta\in\mathcal{A}^{\Theta}$, the classical SDE has a unique strong solution $X_{t}^{\theta}$
  2. 2.
    The boundedness of $\Theta$ ensures uniform estimates:
    $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{s}^{\theta}|^{2}\right]<\infty$
  3. 3.
    The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation
where $\theta_{t}\in\Theta$ is an adapted control process with $\Theta\subset\mathbb{R}$ compact and convex.
4.1 Existence and Uniqueness Theorem
Theorem 4.2 (Classical SDE Solution).
Assume the coefficients satisfy:
  1. 1.
    Lipschitz continuity: For all $x,y\in\mathbb{R}$, $t\in[0,T]$,
    $|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq K|x-y|$ (4.3)
  2. 2.
    Linear growth: For all $x\in\mathbb{R}$, $t\in[0,T]$,
    $|b(t,x)|+|\sigma(t,x)|\leq K(1+|x|)$ (4.4)
  3. 3.
    Bounded controls: $\theta_{t}\in\Theta$ with $\Theta$ compact, so
    $\exists M>0:|\theta_{t}|\leq M\quad\forall t,\omega$ (4.5)
Then for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$ to (4.1).
Proof.
The boundedness of $\theta_{t}$ ensures that the modified diffusion coefficient $\sigma(t,x)\theta_{t}$ satisfies:
$|\sigma(t,x)\theta_{t}-\sigma(t,y)\theta_{t}|$ $\leq M|\sigma(t,x)-\sigma(t,y)|\leq MK|x-y|$
$|\sigma(t,x)\theta_{t}|$ $\leq M|\sigma(t,x)|\leq MK(1+|x|)$
Thus the Lipschitz condition (4.3) and growth condition (4.4) hold for the controlled SDE. Existence and uniqueness follow from standard SDE theory. ∎
4.2 G-SDE Solution via Parameterization
Theorem 4.3 (G-SDE Solution).
Under the same Lipschitz and growth conditions, the G-SDE (4.2) admits a unique solution given by:
$X_{t}=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
where each $X_{t}^{\theta}$ solves the classical SDE (4.1).
Proof.
The solution is constructed in three steps:
  1. 1.
    For each $\theta\in\mathcal{A}^{\Theta}$, solve the classical SDE (4.1)
  2. 2.
    The boundedness of $\Theta$ ensures uniform estimates:
    $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{s}^{\theta}|^{2}\right]\leq C(1+|X_{0}|^{2})e^{Ct}$
  3. 3.
    The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation
The compactness of $\Theta$ is crucial for preserving the solution properties:
[…]
The boundedness of the control set $\Theta$ is essential for:
  • Preserving the Lipschitz and growth conditions under parameterization
  • Ensuring existence and uniqueness of solutions
  • Maintaining moment estimates and path regularity
This framework provides a rigorous bridge from classical SDE theory to G-SDEs, leveraging the bounded control structure to transfer solution properties.
The G-SDE solution is reconstructed from the parameterized solutions:
[…]
The solution inherits properties from the parameterized solutions:
Theorem 4.4 (Properties of Solutions ).
The G-SDE solution satisfies:
  1. 1.
    Moment estimates: $\hat{\mathbb{E}}\left[\sup_{0\leq s\leq t}|X_{s}|^{p}\right]<\infty$ for $p\geq 2$
  2. 2.
    Continuous dependence: If $b_{n}\to b$, $\sigma_{n}\to\sigma$ uniformly, then $X_{t}^{n}\to X_{t}$ in $L_{G}^{2}$
4.3 Geometric Brownian Motion — A Clear Example
Consider the geometric-type G-SDE with volatility uncertainty:
$dX_{t}=\sigma X_{t}dB_{t}+\gamma X_{t}d\langle B\rangle_{t},\quad X_{0}=x_{0}>0$ (4.6)
Through the isomorphism $\Phi$, we transform the G-SDE into a family of classical SDEs:
$\Phi(X_{t})=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
where for each control process $\theta\in\mathcal{A}^{\Theta}$
$dX_{t}^{\theta}=\sigma X_{t}^{\theta}\theta_{t}dW_{t}+\gamma X_{t}^{\theta}\theta_{t}^{2}dt,\quad X_{0}^{\theta}=x_{0}$ (4.7)
Applying the G-Itô formula to $f(x)=\ln x$:
$d(\ln X_{t})$ $=\sigma dB_{t}+\gamma d\langle B\rangle_{t}-\frac{1}{2}\sigma^{2}d\langle B\rangle_{t}$ (4.8)
$=\sigma dB_{t}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)d\langle B\rangle_{t}$ (4.9)
Integrating yields the solution:
$X_{t}=x_{0}\exp\left(\sigma B_{t}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)\langle B\rangle_{t}\right)$ (4.10)
For each fixed control $\theta$, apply classical Itô’s formula:
$d(\ln X_{t}^{\theta})$ $=\sigma\theta_{t}dW_{t}+\gamma\theta_{t}^{2}dt-\frac{1}{2}\sigma^{2}\theta_{t}^{2}dt$ (4.11)
$=\sigma\theta_{t}dW_{t}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)\theta_{t}^{2}dt$ (4.12)
Integrating gives the parameterized solution:
$X_{t}^{\theta}=x_{0}\exp\left(\sigma\int_{0}^{t}\theta_{s}dW_{s}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)\int_{0}^{t}\theta_{s}^{2}ds\right)$ (4.13)
Note that the two are identical under isomorphism.
The G-expectation is obtained by taking supremum over controls:
$\hat{\mathbb{E}}[X_{t}]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[X_{t}^{\theta}]$ (4.14)
Thus the G-expectation is:
$\hat{\mathbb{E}}[X_{t}]=\sup_{\theta\in[\underline{\sigma},\overline{\sigma}]}x_{0}\exp(\gamma\theta^{2}t)=\begin{cases}x_{0}\exp(\gamma\overline{\sigma}^{2}t),&\gamma>0\\ x_{0}\exp(\gamma\underline{\sigma}^{2}t),&\gamma<0\end{cases}$ (4.15)
5 Properties of G-Martingales
In G-expectation theory, martingales exhibit fundamentally different behaviors due to volatility uncertainty. This chapter explains symmetric and asymmetric martingales through the lens of parameterized controls, revealing their economic interpretations as different types of ”fair gambles”.
Under the isomorphism $\Phi$, a G-martingale $M_{t}$ is represented as a family of classical processes:
$\Phi(M_{t})=\{M_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
where each $M_{t}^{\theta}$ is adapted to the filtration generated by $W_{t}^{\theta}$.
5.1 Symmetric Martingales: Fair Gambles in All Scenarios
Definition 5.1 (Symmetric Martingale).
A process $M_{t}$ is a symmetric G-martingale if for every control $\theta\in\mathcal{A}_{s,t}^{\Theta}$, $M_{t}^{\theta}$ is a classical martingale under $\mathbb{P}$:
$\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta}|\mathcal{F}_{s}]=M_{s}^{\theta}\quad\forall s\leq t.$
[…]
5.2 Asymmetric Martingales: Favorable Gambles with Worst-Case Fairness
Definition 5.2 (Asymmetric Martingale).
A process $M_{t}$ is an asymmetric G-martingale if:
  1. 1.
    For all $\theta\in\mathcal{A}_{s,t}^{\Theta}$, $M_{t}^{\theta}$ is a supermartingale under $\mathbb{P}$:
    $\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta}|\mathcal{F}_{s}]\leq M_{s}^{\theta}\quad\forall s\leq t$
  2. 2.
    There exists a worst-case scenario $\theta^{*}\in\Theta$ such that:
    $\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta^{*}}|\mathcal{F}_{s}]=M_{s}^{\theta^{*}}$
5.3 Key Properties
[…]
6 Conclusion
This paper has established a fundamental and rigorous connection between the abstract theory of $G$-expectation and the concrete framework of parameterized stochastic control. By introducing the concept of controlled processes $\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$, we have constructed a complete isomorphism $\Phi$ that maps the nonlinear $G$-expectation space onto a family of classical probability spaces indexed by volatility controls.
The core contribution of this work is threefold:
  1. 1.
    Unified Representation: We demonstrated that $G$-Brownian motion and $G$-SDE solutions admit a natural representation as suprema over parameterized classical processes, providing an intuitive and computationally tractable interpretation of model uncertainty.
  2. 2.
    Algebraic and Analytic Isomorphism: The mapping $\Phi$ was shown to preserve all essential operations – addition, multiplication, function composition, and limit transitions – thereby maintaining the complete mathematical structure while transitioning between nonlinear and classical frameworks.
  3. 3.
    Martingale Classification: Our framework provided a clear distinction between symmetric $G$-martingales (fair in all volatility scenarios) and asymmetric $G$-martingales (favorable but fair only in worst-case scenarios), revealing the fundamental structure of stochastic processes under volatility uncertainty.
The parameterized control perspective developed herein offers more than theoretical elegance; it provides a practical computational methodology for working with $G$-expectations by reducing complex nonlinear problems to families of more manageable classical problems. This approach maintains full mathematical rigor while offering significant advantages for both theoretical analysis and numerical implementation.
Looking forward, this isomorphism establishes a solid foundation for further research in nonlinear stochastic calculus, with particular promise for applications in mathematical finance where volatility uncertainty plays a crucial role in pricing and risk management. The framework naturally extends to more complex stochastic processes and offers a new paradigm for understanding model uncertainty in quantitative economics.

### v7

@@
-Parameterized Controls and Isomorphic Representations of G-Expectations
+A Parametric Methodology for $G$-Expectations and Stochastic Systems under Model Uncertainty
 Abstract
-This paper establishes a fundamental isomorphism between the abstract framework of $G$-expectation and a concrete parameterized family of classical stochastic processes, providing a unified framework for nonlinear stochastic calculus under model uncertainty.
-We demonstrate that $G$-Brownian motion and related stochastic differential equations (SDEs) can be intrinsically represented as controlled processes $\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$, where $\theta$ denotes an admissible control process taking values in a compact convex parameter set $\Theta$ encoding volatility uncertainty.
-The core contribution is the construction of a complete algebraic and analytic isomorphism $\Phi$ that preserves all operations including addition, multiplication, function composition, and limit transitions.
-This isomorphism provides a rigorous mathematical foundation for understanding $G$-martingales, with a distinction between symmetric martingales (fair in all volatility scenarios) and asymmetric martingales (favorable but fair only in the worst-case scenario).
-We develop a comprehensive theory for solving $G$-SDEs through their parameterized representations, establishing existence and uniqueness results under standard Lipschitz and growth conditions.
-The framework reveals how bounded control sets preserve these classical conditions while capturing volatility uncertainty.
-The parameterized control perspective offers both mathematical rigor and computational practicality, enabling the transfer of classical stochastic calculus results to the $G$-expectation framework while maintaining the essential sublinear structure of model uncertainty.
-Key Words: G-Expectation, G-Brownian Motion, Parameterized Controls, Isomorphic Representations, Volatility Uncertainty, Dynamic Programming Principle
+This paper develops a systematic parametric method for analyzing stochastic systems under volatility uncertainty within the $G$-expectation framework.
+Leveraging the dual representation of the $G$-expectation as a supremum over a family of probability measures, we operationalize the correspondence between the $G$-expectation space and a parameterized family of classical stochastic processes.
+Our approach decomposes complex nonlinear analysis into two distinct phases: a “linear implementation phase,” which utilizes classical stochastic analysis under a fixed parameter, and a “consistent estimation phase” across the parameter space.
+We rigorously prove that this parametric mapping preserves fundamental stochastic structures, including the Itô integral and quadratic variation, thus enabling a direct transplantation of classical Itô calculus techniques into the $G$-framework.
+The advantages of this methodology are illustrated through its application to $G$-stochastic differential equations and $G$-backward stochastic differential equations.
+Furthermore, we demonstrate that this method effectively mitigates the inherent accumulation of estimation errors found in traditional stepwise sublinear procedures, yielding significantly more accurate estimates for fundamental quantities in robust stochastic analysis.
+Keywords: $G$-Expectation, Parametric Representation, Model Uncertainty, $G$-SDEs, Uniform Estimation, Financial Applications, Precision Improvement
 E-mail: zhaoguangqian@mail.ustc.edu.cn
 1 Introduction
-The quantification and management of model uncertainty—often referred to as Knightian uncertainty or ambiguity—represents a fundamental challenge across several disciplines, most notably in mathematical finance and decision theory.
-Traditional stochastic calculus, built upon the foundation of a single, precisely known probability measure, is inherently limited in its capacity to address situations where the model itself is imperfectly known.
-This limitation is particularly acute in financial markets, where volatility—the key parameter in option pricing and risk management—is notoriously difficult to estimate and is subject to significant time-varying uncertainty.
-The groundbreaking framework of $G$-expectation, pioneered by Peng , provides a rigorous mathematical foundation for modeling such uncertainty.
-By replacing the classical linear expectation with a sublinear operator, the $G$-framework successfully captures the essence of volatility ambiguity.
-Within this framework, the $G$-Brownian motion $B_{t}$ and related stochastic differential equations ($G$-SDEs) have been defined, with their quadratic variation $\langle B\rangle_{t}$ evolving within a prescribed interval $\underline{\sigma}^{2}t,\overline{\sigma}^{2}t]$, naturally encoding the uncertainty in volatility.
-This leads to a rich theory of nonlinear stochastic calculus .
-A profound insight emerging from this framework is the intrinsic connection between $G$-expectations and stochastic control theory.
-Specifically, the $G$-expectation can be naturally represented as the value function of a stochastic control problem:
-$\hat{\mathbb{E}}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right],$
-where the control process $\theta_{s}$ takes values in a compact set $\Theta$ representing the volatility uncertainty.
-This representation naturally leads to a Hamilton-Jacobi-Bellman (HJB) equation through the Dynamic Programming Principle (DPP):
-$\partial_{t}v+G(D_{x}^{2}v)=0,\quad v(T,x)=\varphi(x),$
-where the nonlinearity $G(A)=\frac{1}{2}\sup_{\gamma\in\Theta}\text{tr}(A\gamma\gamma^{\top})$ encodes the uncertainty set. The DPP manifests infinitesimally as:
-$v(t,x)=\sup_{\theta\in\mathcal{A}_{t,t+h}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[v\left(t+h,x+\int_{t}^{t+h}\theta_{s}dW_{s}\right)\right].$
-Despite this elegant connection, the abstract nature of the $G$-expectation can present a barrier to intuitive understanding and practical application. A central question thus arises:
-Can we establish a more explicit and operational connection between the abstract $G$-expectation space and the concrete framework of parameterized controlled processes?
-This paper provides a definitive affirmative answer by constructing a complete isomorphism between these two domains.
-We demonstrate that the entire $G$-expectation framework admits a rigorous correspondence to a family of classical stochastic processes controlled by volatility parameters.
-The core idea is to represent a $G$-Brownian motion path as a family of classical Itô processes:
-$\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}},\quad\text{where}\quad B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}.$
-The sublinear $G$-expectation is then revealed to be the supremum over all these classical scenarios, providing a stochastic control interpretation that complements the PDE perspective.
-This isomorphic correspondence, denoted by $\Phi$, is shown to preserve the entire algebraic and analytic structure—including addition, multiplication, function composition, and limit transitions—thereby bridging the abstract nonlinear theory with the well-understood classical theory.
-Contributions and Outline: The main contributions of this paper are as follows:
-  • We construct a precise isomorphism $\Phi$ between the $G$-expectation space and a space of parameterized controlled processes, providing a concrete representation that complements the HJB equation characterization (Section 3).
-  • We develop a comprehensive theory for $G$-SDEs, proving existence and uniqueness of solutions through their parameterized representations, thereby connecting the SDE solutions with the dynamic programming framework (Section 4).
-  • Through this isomorphism, we characterize the fundamental structure of $G$-martingales, introducing a distinction between symmetric martingales (fair in all volatility scenarios) and asymmetric martingales (favorable but fair only in the worst-case scenario) (Sections 5 .
-  • We resolve subtle definability difficulties associated with conditional $G$-expectations, clarifying the interplay between supremum operations and conditional expectations under model uncertainty.
-The paper is organized as follows: Section 2 lays the theoretical foundation, formalizing $G$-expectation as a value function and deriving its connection to dynamic programming and nonlinear PDEs.
-Section 3 introduces the fundamental isomorphism, the central concept of this work.
-Sections 4 and 5 delve into the solutions of $G$-SDEs and the properties of $G$-martingales from this new perspective, respectively.
-Finally, Section 6 concludes.
-Our work not only deepens the theoretical understanding of the $G$-framework but also unlocks new computational pathways.
-By translating complex nonlinear problems into families of more tractable classical ones, this parameterized control perspective offers a powerful tool for both theoretical analysis and practical computation in the presence of model uncertainty, while maintaining a natural connection to the HJB and dynamic programming framework.
-2 Theoretical Foundation
+The mathematical treatment of Knightian uncertainty, particularly model ambiguity in volatility, has been fundamentally transformed by the pioneering work of Peng and colleagues through the establishment of the rigorous framework of $G$-expectation theory [1, 3, 4].
+This groundbreaking work provides a powerful, axiomatic foundation for robust financial analysis, enabling the pricing and hedging of derivatives, portfolio optimization, and risk assessment in environments where the underlying volatility is not precisely known but resides within a deterministic set $\underline{\sigma},\overline{\sigma}]$.
+The $G$-expectation framework offers a coherent apparatus with deep connections to stochastic control and fully nonlinear partial differential equations (PDEs), forming a cornerstone for modern quantitative finance under model uncertainty.
+Furthermore, significant advances in related mathematical domains have enriched the landscape of stochastic analysis.
+The framework of quasi-sure analysis developed by Soner, Touzi, and Zhang [9] provides a powerful methodology for aggregating stochastic processes under a non-dominated family of measures, addressing fundamental questions of consistency and robustness.
+In a different direction, Malliavin calculus [42] has emerged as a profound tool for the fine analysis of probability laws on Wiener space, enabling deep sensitivity analysis and the study of regularity for solutions to stochastic systems.
+With continuing research, the axiomatization and intrinsic approach to nonlinear expectation have witnessed significant advances.
+However, as research has matured and delved deeper into the intrinsic properties of the $G$-framework, a subtle yet significant divergence from the vast toolkit of classical stochastic analysis has emerged.
+While the foundational literature—notably Denis, Hu, and Peng [2]—clearly establishes the dual representation of the $G$-expectation as a supremum over a family of linear expectations, the systematic operationalization of this representation for stochastic calculus has remained largely underdeveloped.
+The pathwise dynamics of $G$-processes are often obscured when one works directly with sublinear evaluations or the associated fully nonlinear PDEs.
+This gap has fostered a methodological conservatism in the field.
+There exists a justifiable, yet mathematically constraining, apprehension: the fear that stepping outside the intrinsic, nonlinear formalism might compromise the robustness of the analysis.
+This “mathematical fear” has, in practice, led to an over-reliance on stepwise sublinear procedures.
+While robust by construction, these methods often forcibly decouple the correlation structure between stochastic components by applying sublinear inequalities at intermediate steps.
+As we demonstrate in Section 7, this not only leads to a significant accumulation of estimation errors, yielding excessively conservative and distorted bounds, but also renders the analysis computationally inflexible and inaccessible to those well-versed in classical theory.
+This paper directly addresses this core challenge by constructing a rigorous and systematic parametric methodology that bridges the $G$-expectation space with the classical probability space.
+Our primary contribution is not in identifying the duality—which is known—but in rigorously operationalizing it.
+We explicitly construct a homomorphism $\Phi$ that maps the $G$-expectation space to parameterized families of classical stochastic processes, and we prove that this mapping preserves the essential differential structures, including Itô integrals, quadratic variations, and the dynamics of stochastic differential equations.
+Our framework strategically disentangles the analytical complexity by establishing a clear methodological dichotomy:
+  1. 1.
+    Linear Execution Phase: We demonstrate that the core computational workload—solving SDEs, applying Itô’s lemma, establishing pathwise properties—can be legitimately carried out within the familiar realm of classical linear stochastic analysis for each fixed parameter $\theta$.
+This allows researchers to leverage the entire arsenal of classical probability theory.
+  2. 2.
+    Uniform Estimation Phase: The crucial synthesis into the robust $G$-framework is achieved by a final, well-defined supremum operation, predicated on establishing estimates that hold uniformly across the parameter space $\Theta$.
+This structured separation is transformative.
+It conclusively demonstrates that the “mathematical fear” of losing robustness is unfounded; robustness is guaranteed by the final supremum, provided the intermediate analysis is valid for each scenario.
+By consolidating the analytical challenge solely into the requirement of uniform estimates, our framework dramatically reduces complexity and, as empirically shown in Section 7, significantly improves estimation precision.
+It allows one to enjoy the robustness of the $G$-framework without sacrificing the power and flexibility of classical stochastic calculus.
+The remainder of this paper is organized as follows.
+Section 2 revisits the foundational concepts and establishes the parametric representation.
+Section 3 is devoted to the rigorous construction of the homomorphism $\Phi$ and the proof of its structure-preserving properties.
+Sections 4 and 5 demonstrate the application of this methodology to $G$-SDEs and $G$-BSDEs, respectively.
+Section 6 provides a unified perspective on the homomorphism as a change of probabilistic lenses on the canonical path space.
+Section 7 presents a comparative analysis that quantitatively demonstrates the precision improvement of our approach.
+Section 8 discusses the broader implications of this parametric bridge for future research, and Section 9 concludes.
+2 Fundamental Concepts and Parametric Representation
 2.1 G-Expectation as a Value Function
-The fundamental insight is that G-expectation can be represented as the value function of a stochastic control problem.
+The fundamental insight is that G-expectation can be represented as the value function of a stochastic control problem[1].
 For a G-Brownian motion $B_{t}$ and terminal payoff $\varphi\in C_{b}(\mathbb{R}^{d})$, we have:
@@
 $v(t,x)=\sup_{\theta\in\mathcal{A}^{\Theta}_{t,t+h}}\mathbb{E}_{\mathbb{P}}\left[v\left(t+h,x+\int_{t}^{t+h}\theta_{s}dW_{s}\right)\right]$ (2.5)
-3 The Fundamental Isomorphism: G-Brownian motion as Parameterized Controlled Processes
-The central insight of this work is establishing a natural isomorphism between elements in the G-expectation space and parameterized families of controlled processes. Specifically:
+The central insight of this work is establishing a natural homomorphism between elements in the G-expectation space and parameterized families of controlled processes. Specifically:
   1. 1.
@@
   3. 3.
-    The isomorphism: There exists a bijective correspondence:
+    The homomorphism: There exists a injective correspondence:
     $\begin{array}[]{c}\text{G-expectation space}\\ \Updownarrow\\ \left\{X^{\theta}:\theta\in\Theta\right\}\end{array}$
     where the parameterized family $\{X^{\theta}\}_{\theta\in\Theta}$ is the G-SDE solution
-This isomorphism reveals that:
-  • G-processes are intrinsically controlled trajectories parameterized by volatility uncertainty
-  • The G-expectation framework provides a geometric encoding of model uncertainty
-  • Solutions to G-SDEs emerge as universal families of classical SDE solutions
 2.3 Classical G-Expectation and G-Brownian Motion
@@
 The cornerstone of the theory is the G-normal distribution .
-Let $\xi$ be a random variable.
+Hereafter, we will focus on the one-dimensional case, while the analysis for multi-dimensional situations follows similarly.Let $\xi$ be a random variable.
 It is said to follow a G-normal distribution, denoted $\xi\sim\mathcal{N}(0,[\underline{\sigma}^{2},\overline{\sigma}^{2}])$, if for any $\phi\in C_{b}(\mathbb{R})$, its sublinear expectation is defined by the following nonlinear heat equation:
@@
 $\lim_{t\downarrow 0}\frac{\mathbb{E}[\phi(x+B_{t})]-\phi(x)}{t}=G(\phi^{\prime\prime}(x)).$ (2.10)
-3.2 A Generalized Construction via Random Variables
-This generator governs the evolution of expectations under the G-Brownian motion.
-The drift term is zero, reflecting the martingale property of $B_{t}$, and the diffusion term is given by the supremum over constant parameters.
-This establishes the foundational link: classical G-Brownian motion corresponds to a constant diffusion coefficient uncertainty set $\Theta=[\underline{\sigma},\overline{\sigma}]$.
+2.4 Parameterized Control Representation
+Now we draw upon Peng’s framework to construct the corresponding parameterized control representation.For details of the construction, see [1].
 Definition 2.1 (Construction via Supremum over Linear Expectations).
-Let $(\Omega,\mathcal{F},P)$ be a probability space sufficient to support a standard Brownian motion $\{W_{t}\}_{t\geq 0}$.
-For a given compact, convex parameter set $\Lambda\subset\mathbb{R}^{n}$, we define two fundamental random objects:
-1. Max-Distribution: For a parameter $\mu\in\Lambda$, define the random variable $\eta(\mu)=\mu$. The associated nonlinear expectation $\mathbb{E}[\eta]$ is defined for a test function $\phi$ as:
+Let $(\Omega,\mathcal{F},\mathcal{F}_{t},P)$ be a filtered probability space sufficient to support a standard Brownian motion $\{W_{t}\}_{t\geq 0}$.
+For a given compact, convex parameter set $\Lambda\subset\mathbb{R}^{n}$ and set $\Theta\subset\mathbb{R}^{d\times d}$, we define two fundamental random objects:
+Max-Distribution: For a parameter $\mu\in\Lambda$, define the random variable $\eta(\mu)=\mu$. The associated nonlinear expectation $\mathbb{E}[\eta]$ is defined for a test function $\phi$ as:
 $\mathbb{E}[\phi(\eta)]:=\sup_{\mu\in\Lambda}E_{P}[\phi(\eta(\mu))].$ (2.11)
 This object captures the pure parameter uncertainty in its most extreme form.
-2.
-G-Normal Distribution: For a parameter $\sigma\in\Lambda$, define the random variable $X(\sigma)=\sigma W_{1}$ (which is classically normal distributed, $X(\sigma)\sim N(0,\sigma^{2})$).
-The associated nonlinear expectation $\mathbb{E}[X]$ is defined as:
-$\mathbb{E}[\phi(X)]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(X(\sigma))]=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{1})].$ (3.7)
-This constitutes a G-normal distribution with the uncertainty set $\Lambda$.
-3. Quasi G-Brownian Motion: We define a process $\{B_{t}\}_{t\geq 0}$ by its action on test functions. For any $\phi\in C_{l,Lip}(\mathbb{R})$, its nonlinear expectation is given by:
-$\mathbb{E}[\phi(B_{t})]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t})].$ (3.8)
-Furthermore, for a process of the form $H_{t}=bB_{t}+mt$ (where the drift $m$ is also subject to uncertainty within a convex set), the expectation is defined as:
-$\mathbb{E}[\phi(H_{t})]:=\sup_{(b,m)\in\Lambda^{\prime}}E_{P}[\phi(bW_{t}+mt)],$ (3.9)
-where $\Lambda^{\prime}$ is a compact convex set in $\mathbb{R}^{2}$.
-These constructions are defined on a filtered probability space represented by the quadruple:
-$(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{E})$ (3.10)
-where:
-  • $\Omega$ is the sample space, representing the set of all possible outcomes;
-  • $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$, representing the set of all measurable events;
-  • $(\mathcal{F}_{t})_{t\geq 0}$ is a filtration, i.e., an increasing family of sub-$\sigma$-algebras of $\mathcal{F}$ satisfying $\mathcal{F}_{s}\subseteq\mathcal{F}_{t}$ for all $0\leq s\leq t$, which represents the accumulation of information over time;
-  • $\mathbb{E}$ is the sublinear expectation operator defined in equations (3.6) and (3.7), which generalizes the classical linear expectation to account for model uncertainty.
-Remark 3.1.
-The construction of the ”Quasi G-Brownian Motion” and the related objects herein represents the core methodological approach of this paper.
-It is intended to schematically illustrate the central idea of handling model uncertainty via the supremum over linear expectations.
-It is important to note that the process $\{B_{t}\}$ defined here is for methodological demonstration purposes; while it shares the same one-dimensional distributions with the strict mathematical G-Brownian motion, their joint distributions are entirely different, and thus they are not mathematically equivalent.
-Definition 3.2 (General Construction of Sublinear Expectation via Model Uncertainty).
-Let $(\Omega,\mathcal{F},P)$ be a classical probability space. Let $\Theta$ be a non-empty, compact, and convex parameter set representing the domain of model uncertainty.
-Consider a generalparameter-controlled random variable $X(\theta)$, where $\theta\in\Theta$ is the parameter that dictates the model specification.
-The associated sublinear expectation $\mathbb{E}[\cdot]$ is defined for a test function $\phi$ (typically in the space of Lipschitz continuous functions $C_{l,Lip}(\mathbb{R}^{d})$) by the following operation:
-$\mathbb{E}[\phi(X)]:=\sup_{\theta\in\Theta}E_{P}[\phi(X(\theta))].$ (3.11)
-This means the operator $\mathbb{E}$ acts on the random variable $X$ by a two-step procedure:
-1. For a fixed parameter $\theta\in\Theta$, compute the classical linear expectation $E_{P}$ of the random variable $\phi(X(\theta))$ under the probability measure $P$.
-2. Subsequently, take the supremum of these resulting real numbers over all admissible parameters $\theta$ in the uncertainty set $\Theta$.
-This operator is defined on the sublinear expectation space:
-$(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{E})$ (3.12)
-where:
-  • $\Omega$ is the sample space;
-  • $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$;
-  • $(\mathcal{F}_{t})_{t\geq 0}$ is a filtration representing the information flow over time;
-  • $\mathbb{E}$ is the sublinear expectation operator defined above.
-Remarks:
-  • This general definition subsumes the specific constructions (e.g., Max-Distribution, G-Normal Distribution, Generalized G-Brownian Motion) presented in Definition 3.1 as special cases by choosing the parameter-controlled variable $X(\theta)$ and the uncertainty set $\Theta$ appropriately.
-  • The operator $\mathbb{E}$ defined in this way is sublinear: $\mathbb{E}[X+Y]\leq\mathbb{E}[X]+\mathbb{E}[Y]$ and $\mathbb{E}[\lambda X]=\lambda\mathbb{E}[X]$ for $\lambda\geq 0$.
-It also preserves constants: $\mathbb{E}[c]=c$ for any constant $c$.
-While the parameterized construction provides valuable insight, it is crucial to recognize its fundamental limitation: constant control sets cannot capture the path-dependent nature of true G-Brownian motion.
-Consider the parameterized ”quasi G-Brownian motion” defined by:
-$B_{t}^{\sigma}=\sigma W_{t},\quad\sigma\in\Lambda$ (3.13)
-with nonlinear expectation:
-$\mathbb{E}[\phi(B_{t})]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t})]$ (3.14)
-Key observations:
-  1. 1.
-    One-dimensional distributions match: For fixed $t$,
-    $\text{Law}_{G}(B_{t})\cong\sup_{\sigma\in\Lambda}\text{Law}_{P}(\sigma W_{t})$
-    meaning marginal distributions coincide with true G-Brownian motion.
-  2. 2.
-    Path structure diverges: The joint distributions differ fundamentally:
-    $\mathbb{E}[\phi(B_{t},B_{s})]=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t},\sigma W_{s})]\neq\hat{\mathbb{E}}[\phi(B_{t},B_{s})]$
-    This occurs because constant $\sigma$ cannot capture the time-consistent volatility uncertainty of true G-Brownian motion.
-  3. 3.
-    The Difficulty in Defining Conditional Expectation
-3.3 True G-Brownian
-To overcome these limitations, we must introduce time-varying adaptive controls:
-Definition 3.3 (True G-Brownian Motion via Adaptive Controls).
-Let $\Theta\subset\mathbb{R}^{d\times d}$ be compact convex. Define:
+G-Brownian Motion : Let $\Theta\subset\mathbb{R}^{d\times d}$ be compact convex. Define:
 $B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s},\quad\theta\in\mathcal{A}^{\Theta}$ (2.12)
 with quadratic variation:
-$d\langle B^{\theta}\rangle_{t}=(\theta_{t}\theta_{t}^{\top})dt$ (3.16)
+$d\langle B^{\theta}\rangle_{t}=\theta_{t}^{2}dt$ (2.13)
 The genuine G-Brownian motion is then characterized by:
@@
 where $\Theta\subset\mathbb{R}^{d\times d}$ is a compact convex set representing the volatility uncertainty.
-For the G-Brownian motion $B_{t}$, the conditional expectation is:
-${\mathbb{E}}_{t}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}_{t,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{t}^{T}\theta_{s}dW_{s}\right)\middle|\mathcal{F}_{t}\right]$
-This definition provides the mathematical foundation for conditional expectations in the G-framework, capturing the essence of model uncertainty through the supremum over volatility controls.
+Remark 2.1.
+The compactness of the control set $\Theta$ ensures the existence of a countable dense subset $\Theta_{d}\subset\Theta$, which allows us to approximate the supremum over all controls through this subset:
+$\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{P}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right]=\sup_{\theta\in\mathcal{A}^{\Theta_{d}}}\mathbb{E}_{P}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right].$
+This approximation property stems from the $L^{2}$-continuous dependence on the control parameter and provides analytical convenience in various contexts.
+The introduction of such topological considerations within our parametric framework can facilitate certain analytical techniques, though this aspect is not central to our main development.
 Definition 2.2 (Conditional G-Expectation Operator).
-For a G-Brownian motion $\{B_{t}\}_{t\geq 0}$ and a test function $\varphi\in C_{b.Lip}(\mathbb{R}^{d})$, the conditional G-expectation operator $\hat{\mathbb{E}}_{s}[\cdot]$ is defined as:
-${\mathbb{E}}_{s}[\varphi(B_{t})]=\sup_{\theta\in\mathcal{A}_{s,t}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{s}+\int_{s}^{t}\theta_{u}dW_{u}\right)\middle|\mathcal{F}_{s}\right]$
+For a G-Brownian motion $\{B_{t}\}_{t\geq 0}$ and a test function $\varphi\in C_{b,Lip}(\mathbb{R})$, the conditional G-expectation operator $\hat{\mathbb{E}}_{s}[\cdot]$ is defined as:
+$\mathbb{E}_{s}[\varphi(B_{t})]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{s}+\int_{s}^{t}\theta_{u}dW_{u}\right)\Big|\mathcal{F}_{s}\right]$
 where:
-  • $\mathcal{A}_{s,t}^{\Theta}$ denotes admissible control processes on $[s,t]$
-  • $\theta_{u}\in\Theta=[\underline{\sigma},\overline{\sigma}]$ represents volatility controls
-  • $W_{u}$ is a standard Brownian motion under $\mathbb{P}$
+  • $\mathcal{A}_{s,T}^{\Theta}$ denotes admissible control processes on $[s,T]$,
+  • $\theta_{u}\in\Theta=[\underline{\sigma},\overline{\sigma}]$ represents volatility controls,
+  • $W_{u}$ is a standard Brownian motion under $\mathbb{P}$.
 Definition 2.3 (Conditional G-Expectation for Cylinder Functions).
 For any cylinder function of the form:
-$\xi=\varphi(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}}),\quad\varphi\in C_{b.Lip}(\mathbb{R}^{d\times n})$
+$\xi=\varphi(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}}),\quad\varphi\in C_{b,Lip}(\mathbb{R}^{n})$
 with $0\leq t_{1}<t_{2}<\cdots<t_{n}\leq T$, the conditional G-expectation is defined as:
-${\mathbb{E}}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\middle|\mathcal{F}_{s}\right]$
-where $B_{t}^{\theta}=\int_{0}^{t}\theta_{u}dW_{u}$ is the controlled process.
-This definition coincides with the classical G-expectation definition when $s=0$:
-${\mathbb{E}}[\xi]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right]$
+$\mathbb{E}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}}\right)\Big|\mathcal{F}_{s}\right]$
+Note: In the above, the control process $\theta$ affects the increments after time $s$. For increments before time $s$, they are fixed by the observation.
 Theorem 2.4 (Conditional Expectation Simplification).
@@
 $\mathbb{E}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(x_{1},x_{2},\cdots,x_{k-1},B_{t_{k}}^{\theta}-B_{t_{k-1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right]\Bigg|_{x_{i}=B_{t_{i}}}$
-where for $t_{k-1}\leq s$$t_{k}>s$, we treat $B_{t_{i}}$ as constants $x_{i}$ in the expectation.
-Note that our definition of conditional expectation here is completely consistent with the classical definition of G-conditional expectation under the definition of cylindrical Brownian motion, so we can similarly extend it to its completion space.
+where for $t_{k-1}\leq s<t_{k}$, we treat $B_{t_{i}}$ for $i=1,\dots,k-1$ as constants $x_{i}$ in the expectation, and for $t\geq s$, we define $B_{t}^{\theta}=B_{s}+\int_{s}^{t}\theta_{u}dW_{u}$.
+This definition coincides with the classical G-expectation definition when $s=0$:
+$\mathbb{E}[\xi]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right].$
+Note that our definition of conditional expectation is completely consistent with the classical definition of G-conditional expectation under cylindrical Brownian motion, allowing for natural extension to the completion space.
+This framework extends seamlessly to stopping times, where conditioning occurs on the information up to the stopping time $\tau$.
+Furthermore, the elements preceding the conditional expectation, as members of the G-expectation space, naturally correspond to elements in the parameterized family.
+For notational simplicity, we have omitted explicit parameterization in these expressions, though the underlying parametric structure remains intact throughout our framework.
 Theorem 2.5 (Fundamental Properties).
@@
 This ensures the time consistency of the G-expectation framework.
-3.4 The Complete Isomorphism
-To establish a full isomorphism, we must introduce time-varying adaptive controls.
-3.4.1 Adaptive Control Representation
-Definition 3.9 (True G-Brownian Motion via Adaptive Controls).
-Let $\Theta\subset\mathbb{R}$ be compact convex. Define:
-$B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s},\quad\theta\in\mathcal{A}^{\Theta}$ (3.18)
-with quadratic variation:
-$d\langle B^{\theta}\rangle_{t}=\theta_{t}^{2}dt$ (3.19)
-The genuine G-Brownian motion is characterized by:
-$\hat{\mathbb{E}}[\phi(B_{t})]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[\phi(B_{t}^{\theta})]$ (3.20)
-3.4.2 Algebraic Structure Preservation
-The isomorphism preserves all algebraic operations:
-Theorem 3.10 (Algebraic Isomorphism).
-The mapping $\Phi:X\mapsto\{X^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ preserves:
-  1. 1.
-    Addition: $\Phi(X+Y)=\Phi(X)+\Phi(Y)$
-  2. 2.
-    Scalar Multiplication: $\Phi(\lambda X)=\lambda\Phi(X)$
-  3. 3.
-    Multiplication: $\Phi(X\cdot Y)=\Phi(X)\cdot\Phi(Y)$
-  4. 4.
-    Function Composition: $\Phi(\phi(X))=\phi(\Phi(X))$
-The preservation follows directly from the pathwise definition:
-$(X+Y)^{\theta}$ $=X^{\theta}+Y^{\theta}$
-$(\lambda X)^{\theta}$ $=\lambda X^{\theta}$
-$(X\cdot Y)^{\theta}$ $=X^{\theta}\cdot Y^{\theta}$
-$\phi(X)^{\theta}$ $=\phi(X^{\theta})$
-Fourth, we define the multivariate composition in a similar manner.
-Building upon the correspondence established via G-Brownian motion, we derive a corresponding relationship in the space spanned by these two elements, while ensuring that the mapping constitutes a bijection.
-3.4.3 Analytic Structure Preservation
-The isomorphism also preserves analytic properties:
-Theorem 3.11 (Analytic Isomorphism).
+3 The homomorphism
+This chapter is devoted to the rigorous construction of the fundamental homomorphism $\Phi$ between the space of random variables under the G-expectation and the space of parameterized families of classical random variables.
+The central goal is to demonstrate that this mapping not only exists but also preserves the essential algebraic structure of the spaces involved.
+The construction proceeds in a canonical manner: we first define $\Phi$ explicitly on a dense subset of cylinder functions, establish its key properties there, and then extend it continuously to the entire space.
+The preservation of addition, scalar multiplication, and multiplication under this extension is then proven.
+3.1 Notation and Preliminaries
+Let $(\Omega,\mathcal{F},P)$ be a complete probability space supporting a standard Brownian motion $\{W_{t}\}_{t\geq 0}$, and let $\{\mathcal{F}_{t}\}$ be its natural filtration, augmented to satisfy the usual conditions.
+Let $\Theta\subset\mathbb{R}$ be a compact convex set representing the volatility uncertainty.
+The set of admissible controls is defined as:
+$\mathcal{A}^{\Theta}:=\left\{\theta:[0,\infty)\times\Omega\to\Theta\mid\theta\text{ is }\{\mathcal{F}_{t}\}\text{-progressively measurable and }E_{P}\left[\int_{0}^{T}|\theta_{s}|^{2}ds\right]<\infty\ \forall T>0\right\}.$
+For any $\theta\in\mathcal{A}^{\Theta}$, the controlled process is defined by:
+$B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}.$
+The G-expectation space $L_{G}^{p}(\Omega_{T})$ is the completion of the space of cylinder functions under the norm $\|X\|_{L_{G}^{p}}=\hat{E}[|X|^{p}]^{1/p}$, where $\hat{E}$ denotes the G-expectation operator.
+A cylinder function is a random variable of the form:
+$\xi=\varphi\left(B_{t_{1}\wedge T},B_{t_{2}\wedge T},\cdots,B_{t_{n}\wedge T}\right),\quad\forall n\in\mathbb{N},t_{1},\ldots,t_{n}\in[0,\infty),\varphi\in C_{l.Lip}(\mathbb{R}^{n})$
+where $0\leq t_{1}<t_{2}<\ldots<t_{n}\leq T$ and $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is a local Lipschitz continuous function.
+The collection of all such cylinder functions is denoted by $\mathcal{C}_{T}$.
+Lemma 3.1 (Density of Cylinder Functions).
+The set $\mathcal{C}_{T}$ is dense in $L_{G}^{p}(\Omega_{T})$ for any $p\geq 1$.
+Proof.
+This is a foundational result in the theory of G-expectation. See, for example, [1]
+3.2 Construction of the homomorphism $\Phi$
+We now construct the mapping $\Phi:L_{G}^{p}(\Omega_{T})\to{P}$, where P{P} denotes the space of families of random variables $\{X^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ indexed by the control, with appropriate integrability conditions.
+To provide rigorous topological foundations for the homomorphism $\Phi$, we define the target space as a Banach space of parameterized random variables.
+Definition 3.2 (Uniform $L^{p}$-Norm).
+Let $\mathbf{X}=\{X^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ be a family of random variables with $X^{\theta}\in L^{p}(P)$ for each $\theta$. The uniform norm is defined as:
+$\|\mathbf{X}\|_{\mathcal{P}_{p}}:=\sup_{\theta\in\mathcal{A}^{\Theta}}\|X^{\theta}\|_{L^{p}(\mathbb{P})}.$
+Definition 3.3 (Parameterized Cylinder Functions).
+The space of parameterized cylinder functions is:
+$\mathcal{C}_{par}=\left\{\left\{\varphi(B_{t_{1}}^{\theta},\ldots,B_{t_{n}}^{\theta})\right\}_{\theta\in\mathcal{A}^{\Theta}}:\varphi\in C_{l.Lip}(\mathbb{R}^{n}),\ t_{i}\geq 0\right\}.$
+Definition 3.4 (Target Space $\mathcal{P}_{p}$).
+The target space is the completion:
+$\mathcal{P}_{p}:=\overline{\mathcal{C}_{par}}^{\|\cdot\|_{\mathcal{P}_{p}}},$
+which forms a Banach space under the uniform norm.
+Remark 3.1.
+With this topology, the convergence condition in Definition 3.4 becomes standard Banach space convergence:
+$\|\Phi(X)-\Phi(\xi_{n})\|_{\mathcal{P}_{p}}\to 0,$
+ensuring the well-definedness of $\Phi$ on $L_{G}^{p}$.
+Definition 3.5 (Definition on Cylinder Functions).
+For any cylinder function $\xi=\varphi(B_{t_{1}\wedge T},\ldots,B_{t_{n}\wedge T})\in\mathcal{C}_{T}$, we define:
+$\Phi(\xi):=\left\{\xi^{\theta}\right\}_{\theta\in\mathcal{A}^{\Theta}},\quad\text{where}\quad\xi^{\theta}:=\varphi(B_{t_{1}\wedge T}^{\theta},\ldots,B_{t_{n}\wedge T}^{\theta}).$
+A crucial property of this definition is the preservation of the G-norm, which follows from the definition of G-expectation as a supremum over controls.
+Lemma 3.6 (Isometry on Cylinder Functions).
+For any cylinder function $\xi\in\mathcal{C}_{T}$, we have:
+$\hat{E}[|\xi|^{p}]=\sup_{\theta\in\mathcal{A}^{\Theta}}E_{P}[|\xi^{\theta}|^{p}].$
+Consequently, $\|\xi\|_{L_{G}^{p}}=\sup_{\theta}\|\xi^{\theta}\|_{L^{p}(P)}$.
+Proof.
+This follows directly from the representation of the G-expectation for cylinder functions. See[1]. ∎
+This isometry property allows us to extend $\Phi$ to the entire space $L_{G}^{p}(\Omega_{T})$ by continuity.
+Definition 3.7 (Extension by Continuity).
+For any $X\in L_{G}^{p}(\Omega_{T})$, let $\{\xi_{n}\}\subset\mathcal{C}_{T}$ be a sequence such that $\|X-\xi_{n}\|_{L_{G}^{p}}\to 0$. We define:
+$\Phi(X):=\lim_{n\to\infty}\Phi(\xi_{n}),$
+where the limit is taken in the sense that:
+$\sup_{\theta\in\mathcal{A}^{\Theta}}\|X^{\theta}-\xi_{n}^{\theta}\|_{L^{p}(P)}\to 0\quad\text{as}\quad n\to\infty.$
+3.3 Preservation of Algebraic Operations
+We now prove that the homomorphism $\Phi$ preserves the algebraic operations of addition, scalar multiplication, and multiplication.
+Theorem 3.8 (Preservation of Addition).
+For any $X,Y\in L_{G}^{p}(\Omega_{T})$, we have:
+$\Phi(X+Y)=\Phi(X)+\Phi(Y).$
+Proof.
+Let $\{\xi_{n}\},\{\eta_{n}\}\subset\mathcal{C}_{T}$ be sequences approximating $X$ and $Y$, respectively. Then $\{\xi_{n}+\eta_{n}\}$ is a sequence of cylinder functions approximating $X+Y$.
+By the definition on cylinder functions, for each $n$:
+$\Phi(\xi_{n}+\eta_{n})=\{(\xi_{n}+\eta_{n})^{\theta}\}=\{\xi_{n}^{\theta}+\eta_{n}^{\theta}\}=\Phi(\xi_{n})+\Phi(\eta_{n}).$
+Taking the limit as $n\to\infty$ on both sides, and using the continuity of $\Phi$ (which follows from its isometry property), we obtain:
+$\Phi(X+Y)=\lim_{n\to\infty}\Phi(\xi_{n}+\eta_{n})=\lim_{n\to\infty}[\Phi(\xi_{n})+\Phi(\eta_{n})]=\Phi(X)+\Phi(Y).$
+Note that the last equality holds because the limit of a sum is the sum of the limits in the topology of uniform convergence in $L^{p}(P)$ over $\theta$. ∎
+Theorem 3.9 (Preservation of Scalar Multiplication).
+For any $X\in L_{G}^{p}(\Omega_{T})$ and $\lambda\in\mathbb{R}$, we have:
+$\Phi(\lambda X)=\lambda\Phi(X).$
+Proof.
+The proof is analogous. Let $\{\xi_{n}\}\subset\mathcal{C}_{T}$ approximate $X$. Then $\{\lambda\xi_{n}\}$ approximates $\lambda X$. For each $n$:
+$\Phi(\lambda\xi_{n})=\{(\lambda\xi_{n})^{\theta}\}=\{\lambda\xi_{n}^{\theta}\}=\lambda\Phi(\xi_{n}).$
+Taking limits yields the result. ∎
+Theorem 3.10 (Preservation of Multiplication).
+For any $X,Y\in L_{G}^{p}(\Omega_{T})$, where $p\geq 2$,we have:
+$\Phi(X\cdot Y)=\Phi(X)\cdot\Phi(Y).$
+Proof.
+Let $\{\xi_{n}\},\{\eta_{n}\}\subset\mathcal{C}_{T}$ approximate $X$ and $Y$, respectively.
+Note that the product $\xi_{n}\eta_{n}$ is also a cylinder function (by considering a function that multiplies the corresponding components).
+By the definition on cylinder functions:
+$\Phi(\xi_{n}\eta_{n})=\{(\xi_{n}\eta_{n})^{\theta}\}=\{\xi_{n}^{\theta}\eta_{n}^{\theta}\}=\Phi(\xi_{n})\cdot\Phi(\eta_{n}).$
+We now show that $\{\xi_{n}\eta_{n}\}$ approximates $X\cdot Y$ in the parameterized framework. For each $\theta\in\mathcal{A}^{\Theta}$, applying the classical Hölder inequality:
+$\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}\eta_{n}^{\theta}-X^{\theta}Y^{\theta}|^{p/2}]\leq\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}|^{p}]^{1/2}\mathbb{E}_{\mathbb{P}}[|\eta_{n}^{\theta}-Y^{\theta}|^{p}]^{1/2}+\mathbb{E}_{\mathbb{P}}[|Y^{\theta}|^{p}]^{1/2}\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}-X^{\theta}|^{p}]^{1/2}.$
+Since $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}-X^{\theta}|^{p}]\to 0$ and $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[|\eta_{n}^{\theta}-Y^{\theta}|^{p}]\to 0$, and the sequences $\{\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}|^{p}]\}$ and $\{\mathbb{E}_{\mathbb{P}}[|Y^{\theta}|^{p}]\}$ are uniformly bounded in $\theta$, we obtain:
+$\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}\eta_{n}^{\theta}-X^{\theta}Y^{\theta}|^{p/2}]\to 0.$
+This uniform convergence in $L^{p/2}(\mathbb{P})$ over $\theta$ implies:
+$\Phi(X\cdot Y)=\lim_{n\to\infty}\Phi(\xi_{n}\eta_{n})=\lim_{n\to\infty}[\Phi(\xi_{n})\cdot\Phi(\eta_{n})]=\Phi(X)\cdot\Phi(Y).$
+Remark 3.2.
+The preservation of continuous function composition $\Phi(\phi(X))=\phi(\Phi(X))$ relies on the existence of a common subsequence $\{\xi_{n_{k}}\}$ of cylinder functions with specific convergence properties.
+In the original G-expectation space, this subsequence converges quasi-surely to $X$, defining an $L^{p}_{G}$ equivalence class.
+Crucially, in the parameterized space, for each $\theta\in\mathcal{A}^{\Theta}$, the same subsequence converges almost everywhere to $X^{\theta}$ under the probability measure $\mathbb{P}$, while also maintaining $L^{p}$-norm convergence uniformly in $\theta$.
+3.4 Analytic Structure Preservation
+The homomorphism also preserves analytic properties:
+Theorem 3.11 (Analytic homomorphism).
 The mapping $\Phi$ preserves:
@@
   2. 2.
-    Conditional Expectation: $\hat{\mathbb{E}}_{t}[X]=\sup_{\theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}|\mathcal{F}_{t}]$
+    Conditional Expectation: $\hat{\mathbb{E}}_{t}[X]=\sup\limits_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
   3. 3.
     Convergence: $X_{n}\to X$ in $L_{G}^{p}$ iff $X_{n}^{\theta}\to X^{\theta}$ in $L^{p}$ uniformly over $\theta$
-3.5 Stochastic Calculus Equivalence
-The isomorphism extends to differential operations:
+The homomorphism extends to differential operations:
 3.4.1 Itô Calculus Preservation
-Theorem 3.12 (Itô Calculus Isomorphism).
+Theorem 3.12 (Itô Calculus homomorphism).
 For G-Itô processes:
@@
   1. 1.
-    $\Phi(dB_{t})=\{\theta_{t}dW_{t}\}$
+    $\Phi\left(\int_{0}^{t}dB_{s}\right)=\left\{\int_{0}^{t}\theta_{s}dW_{s}\right\}$
   2. 2.
@@
     $\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}$
-3.5.2 Itô Calculus Isomorphism (Integral Form)
+3.4.2 Itô Calculus homomorphism (Integral Form)
 Proof.
@@
 Part 1: $\Phi\left(\int_{0}^{t}dB_{s}\right)=\left\{\int_{0}^{t}\theta_{s}dW_{s}\right\}$
-By definition of the isomorphism:
+By definition of the homomorphism
 $\Phi\left(\int_{0}^{t}dB_{s}\right)=\Phi(B_{t})=\left\{B_{t}^{\theta}\right\}_{\theta\in\mathcal{A}^{\Theta}}=\left\{\int_{0}^{t}\theta_{s}dW_{s}\right\}_{\theta\in\mathcal{A}^{\Theta}}$
@@
 $\Phi\left(\int_{0}^{t}\partial_{x}f(B_{s})dB_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s})d\langle B\rangle_{s}\right)=\left\{\int_{0}^{t}\partial_{x}f(B_{s}^{\theta})\theta_{s}dW_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s}^{\theta})\theta_{s}^{2}ds\right\}_{\theta\in\mathcal{A}^{\Theta}}$
-This confirms the isomorphism preserves the integral structure of Itô calculus in L²-sense. ∎
-3.6 Visualizing the Isomorphism: From G-SDE to Parameterized Processes
+This confirms the homomorphism preserves the integral structure of Itô calculus in L²-sense. ∎
+3.5 Schematic of G-SDE Analysis Steps under the Parametric Framework
 […]
-The isomorphism $\Phi$ establishes a complete correspondence between the G-expectation space and the parameterized classical space:
-with the mapping defined as:
-$\Phi(X)=\{X^{\theta}\}_{\theta\in\Theta}$
+Remark 3.3.
+In the parametric framework for solving G-SDEs, beyond performing conventional operations, it is crucial to ensure that all convergences hold uniformly in the control parameter with respect to the $L^{p}$-norm.
+This uniform convergence in $\theta$ is precisely what guarantees the well-definedness of the inverse mapping $\Phi^{-1}$, allowing the reconstruction of a single $G$-random variable from the parameterized family $\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$.
+Essentially, our approach consolidates the analytical complexity solely into the final stage of uniform estimation, rather than dispersing it throughout the computational process.
+These uniformity conditions are not only necessary but also economically minimal—they represent the least restrictive requirements to preserve the structural integrity of the homomorphism.
+This underscores the operational optimality of our parametric representation: it leverages classical tools to their fullest extent while seamlessly bridging to the $G$-expectation framework through a single, coherent consistency analysis.
+By adhering to this principle, the parametric representation maintains its mathematical meaning and provides a robust foundation for aggregating the solution family into a universal process on the path space.
 3.6 Algebraic and Analytic Structure Preservation
-The isomorphism preserves all algebraic and analytic operations:
+The homomorphism preserves all algebraic and analytic operations:
 Operation G-Expectation Space Parameterized Space
@@
 Quadratic Variation $\langle B\rangle_{t}$ $\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}$
-The sublinear expectation operators are related through the isomorphism:
+The sublinear expectation operators are related through the injective homomorphism as follows:
 G-Expectation Space Parameterized Space
-$\hat{\mathbb{E}}[X]$ $\sup\limits_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
-$\hat{\mathbb{E}}_{t}[X]$ $\sup\limits_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}\mid\mathcal{F}_{t}]$
+$\hat{\mathbb{E}}[X]$ $\sup\limits_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
+$\hat{\mathbb{E}}_{t}[X]$ $\sup\limits_{\theta\in\mathcal{A}_{t,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}[X^{\theta}\mid\mathcal{F}_{t}]$
+Note that the control sets differ between the unconditional and conditional expectations: the unconditional expectation uses controls over the entire interval $[0,T]$, while the conditional expectation uses controls only over the remaining interval $[t,T]$.
   1. 1.
@@
   3. 3.
-    Analytic Continuity: The isomorphism preserves convergence properties:
+    Analytic Continuity: The homomorphism preserves convergence properties:
     $X_{n}\to X\text{ in }{L^{p}_{\mathbb{G}}}\iff X_{n}^{\theta}\to X^{\theta}\text{ in }L^{p}\text{ uniformly over }\theta$
-This isomorphism provides:
-  • A computational bridge from nonlinear to linear stochastic calculus
-  • A theoretical foundation for model uncertainty quantification
-  • An operational framework for solving G-SDEs through parameterized classical SDEs
 4 How to Solve G-Stochastic Differential Equations
 Consider a classical SDE parameterized by $\theta\in\Theta$:
-$dX_{t}^{\theta}=b(t,X_{t}^{\theta})dt+\sigma(t,X_{t}^{\theta})\theta_{t}dW_{t}$ (4.1)
+$dX_{t}^{\theta}=b(t,X_{t}^{\theta})dt+h(t,X_{t}^{\theta})\theta_{t}^{2}dt+\sigma(t,X_{t}^{\theta})\theta_{t}dW_{t}$ (4.1)
 where:
@@
   • $\Theta\subset\mathbb{R}$ is compact and convex (bounded controls)
-Under standard Lipschitz and growth conditions on $b$ and $\sigma$, for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$.
 We reformulate the classical SDE as a G-SDE driven by G-Brownian motion $B_{t}$:
 The classical SDE (4.1) is equivalent to the following G-SDE:
-$dX_{t}=b(t,X_{t})dt+\sigma(t,X_{t})dB_{t}$ (4.2)
-where $B_{t}$ is a G-Brownian motion
-Theorem 4.1 (Existence and Uniqueness for G-SDE).
-Under bounded controls $\theta_{t}\in\Theta$ and standard Lipschitz conditions:
-$|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|$ $\leq K|x-y|$
-$|b(t,x)|^{2}+|\sigma(t,x)|^{2}$ $\leq K(1+|x|^{2})$
-the G-SDE (4.2) admits a unique solution in the quasi-sure sense:
-$X_{t}=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
-Proof.
-The proof follows from:
-  1. 1.
-    For each fixed $\theta\in\mathcal{A}^{\Theta}$, the classical SDE has a unique strong solution $X_{t}^{\theta}$
-  2. 2.
-    The boundedness of $\Theta$ ensures uniform estimates:
-    $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{s}^{\theta}|^{2}\right]<\infty$
-  3. 3.
-    The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation
-where $\theta_{t}\in\Theta$ is an adapted control process with $\Theta\subset\mathbb{R}$ compact and convex.
-4.1 Existence and Uniqueness Theorem
-Theorem 4.2 (Classical SDE Solution).
+$dX_{t}=b(t,X_{t})dt+h(t,X_{t})d\langle B\rangle_{t}+\sigma(t,X_{t})dB_{t}$ (4.2)
+where $B_{t}$ is a G-Brownian motion.
+4.1 Existence and Uniqueness for the Parameterized Family
+Theorem 4.1 (Classical SDE Solution with Quadratic Variation Term).
 Assume the coefficients satisfy:
@@
     Lipschitz continuity: For all $x,y\in\mathbb{R}$, $t\in[0,T]$,
-    $|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq K|x-y|$ (4.3)
+    $|b(t,x)-b(t,y)|+|h(t,x)-h(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq K|x-y|$ (4.3)
   2. 2.
     Linear growth: For all $x\in\mathbb{R}$, $t\in[0,T]$,
-    $|b(t,x)|+|\sigma(t,x)|\leq K(1+|x|)$ (4.4)
+    $|b(t,x)|+|h(t,x)|+|\sigma(t,x)|\leq K(1+|x|)$ (4.4)
   3. 3.
@@
     $\exists M>0:|\theta_{t}|\leq M\quad\forall t,\omega$ (4.5)
-Then for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$ to (4.1).
+Then for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$ to (4.1), and the solution family satisfies the uniform $L^{2}$-estimate:
+$\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{\theta}|^{2}\right]\leq C(1+|X_{0}|^{2})e^{CT}$ (4.6)
+for some constant $C>0$ depending only on $K$ and $M$.
 Proof.
-The boundedness of $\theta_{t}$ ensures that the modified diffusion coefficient $\sigma(t,x)\theta_{t}$ satisfies:
-$|\sigma(t,x)\theta_{t}-\sigma(t,y)\theta_{t}|$ $\leq M|\sigma(t,x)-\sigma(t,y)|\leq MK|x-y|$
-$|\sigma(t,x)\theta_{t}|$ $\leq M|\sigma(t,x)|\leq MK(1+|x|)$
-Thus the Lipschitz condition (4.3) and growth condition (4.4) hold for the controlled SDE. Existence and uniqueness follow from standard SDE theory. ∎
-4.2 G-SDE Solution via Parameterization
-Theorem 4.3 (G-SDE Solution).
+We prove this via Picard iteration with uniform estimates. Define the Picard iteration scheme:
+$X_{t}^{\theta,(0)}=X_{0},\quad X_{t}^{\theta,(n+1)}=X_{0}+\int_{0}^{t}b(s,X_{s}^{\theta,(n)})ds+\int_{0}^{t}h(s,X_{s}^{\theta,(n)})\theta_{s}^{2}ds+\int_{0}^{t}\sigma(s,X_{s}^{\theta,(n)})\theta_{s}dW_{s}.$
+Step 1: Uniform $L^{2}$-bounds for iterates. For fixed $\theta$ and $n$, by Itô’s isometry and the growth conditions:
+$\mathbb{E}[|X_{t}^{\theta,(n+1)}|^{2}]$ $\leq 4|X_{0}|^{2}+4\mathbb{E}\left[\left|\int_{0}^{t}b(s,X_{s}^{\theta,(n)})ds\right|^{2}\right]+4\mathbb{E}\left[\left|\int_{0}^{t}h(s,X_{s}^{\theta,(n)})\theta_{s}^{2}ds\right|^{2}\right]$
+$\quad+4\mathbb{E}\left[\left|\int_{0}^{t}\sigma(s,X_{s}^{\theta,(n)})\theta_{s}dW_{s}\right|^{2}\right]$
+$\leq 4|X_{0}|^{2}+4t\mathbb{E}\left[\int_{0}^{t}|b(s,X_{s}^{\theta,(n)})|^{2}ds\right]+4t\mathbb{E}\left[\int_{0}^{t}|h(s,X_{s}^{\theta,(n)})\theta_{s}^{2}|^{2}ds\right]$
+$\quad+4\mathbb{E}\left[\int_{0}^{t}|\sigma(s,X_{s}^{\theta,(n)})\theta_{s}|^{2}ds\right]$
+$\leq 4|X_{0}|^{2}+4tK^{2}\mathbb{E}\left[\int_{0}^{t}(1+|X_{s}^{\theta,(n)}|)^{2}ds\right]+4tM^{4}K^{2}\mathbb{E}\left[\int_{0}^{t}(1+|X_{s}^{\theta,(n)}|)^{2}ds\right]$
+$\quad+4M^{2}K^{2}\mathbb{E}\left[\int_{0}^{t}(1+|X_{s}^{\theta,(n)}|)^{2}ds\right]$
+$=4|X_{0}|^{2}+4\left[t(1+M^{4})K^{2}+M^{2}K^{2}\right]\mathbb{E}\left[\int_{0}^{t}(1+|X_{s}^{\theta,(n)}|)^{2}ds\right]$
+$\leq 4|X_{0}|^{2}+8\left[T(1+M^{4})K^{2}+M^{2}K^{2}\right]\left(t+\int_{0}^{t}\mathbb{E}[|X_{s}^{\theta,(n)}|^{2}]ds\right).$
+Let $C_{1}=8\left[T(1+M^{4})K^{2}+M^{2}K^{2}\right]$. Then:
+$\mathbb{E}[|X_{t}^{\theta,(n+1)}|^{2}]\leq 4|X_{0}|^{2}+C_{1}\left(t+\int_{0}^{t}\mathbb{E}[|X_{s}^{\theta,(n)}|^{2}]ds\right).$
+By Gronwall’s inequality, there exists $C_{2}>0$ such that:
+$\sup_{n\geq 0}\mathbb{E}[|X_{t}^{\theta,(n)}|^{2}]\leq C_{2}(1+|X_{0}|^{2})e^{C_{2}t}.$
+Step 2: Uniform convergence of iterates. For $n\geq 1$, consider:
+$\mathbb{E}[|X_{t}^{\theta,(n+1)}-X_{t}^{\theta,(n)}|^{2}]$
+$\leq 3\mathbb{E}\left[\left|\int_{0}^{t}(b(s,X_{s}^{\theta,(n)})-b(s,X_{s}^{\theta,(n-1)}))ds\right|^{2}\right]$
+$\quad+3\mathbb{E}\left[\left|\int_{0}^{t}(h(s,X_{s}^{\theta,(n)})-h(s,X_{s}^{\theta,(n-1)}))\theta_{s}^{2}ds\right|^{2}\right]$
+$\quad+3\mathbb{E}\left[\left|\int_{0}^{t}(\sigma(s,X_{s}^{\theta,(n)})-\sigma(s,X_{s}^{\theta,(n-1)}))\theta_{s}dW_{s}\right|^{2}\right]$
+$\leq 3\left[t+tM^{4}+M^{2}\right]K^{2}\int_{0}^{t}\mathbb{E}[|X_{s}^{\theta,(n)}-X_{s}^{\theta,(n-1)}|^{2}]ds.$
+Let $C_{3}=3(T+TM^{4}+M^{2})K^{2}$ and define $\phi_{n}(t)=\mathbb{E}[|X_{t}^{\theta,(n+1)}-X_{t}^{\theta,(n)}|^{2}]$. Then:
+$\phi_{n}(t)\leq C_{3}\int_{0}^{t}\phi_{n-1}(s)ds.$
+By induction, $\phi_{n}(t)\leq\frac{(C_{3}t)^{n}}{n!}\phi_{0}(T)$.
+Since $\sum_{n=0}^{\infty}\sqrt{\frac{(C_{3}T)^{n}}{n!}}<\infty$, the sequence $\{X_{t}^{\theta,(n)}\}$ converges uniformly in $L^{2}$ to some limit $X_{t}^{\theta}$.
+Step 3: Verification that limit solves SDE. The limit $X_{t}^{\theta}$ satisfies:
+$X_{t}^{\theta}=X_{0}+\int_{0}^{t}b(s,X_{s}^{\theta})ds+\int_{0}^{t}h(s,X_{s}^{\theta})\theta_{s}^{2}ds+\int_{0}^{t}\sigma(s,X_{s}^{\theta})\theta_{s}dW_{s}\quad\text{in }L^{2}.$
+The stochastic integral is well-defined since:
+$\mathbb{E}\left[\int_{0}^{T}|\sigma(s,X_{s}^{\theta})\theta_{s}|^{2}ds\right]\leq M^{2}K^{2}\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]<\infty.$
+Step 4: Uniform $L^{2}$-estimate for solutions. Taking supremum and expectation in the SDE:
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{\theta}|^{2}\right]$ $\leq 4|X_{0}|^{2}+4\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\int_{0}^{t}b(s,X_{s}^{\theta})ds\right|^{2}\right]$
+$\quad+4\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\int_{0}^{t}h(s,X_{s}^{\theta})\theta_{s}^{2}ds\right|^{2}\right]$
+$\quad+4\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\int_{0}^{t}\sigma(s,X_{s}^{\theta})\theta_{s}dW_{s}\right|^{2}\right]$
+$\leq 4|X_{0}|^{2}+4T\mathbb{E}\left[\int_{0}^{T}|b(s,X_{s}^{\theta})|^{2}ds\right]+4T\mathbb{E}\left[\int_{0}^{T}|h(s,X_{s}^{\theta})\theta_{s}^{2}|^{2}ds\right]$
+$\quad+16\mathbb{E}\left[\int_{0}^{T}|\sigma(s,X_{s}^{\theta})\theta_{s}|^{2}ds\right]$
+$\leq 4|X_{0}|^{2}+4TK^{2}\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]+4TM^{4}K^{2}\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]$
+$\quad+16M^{2}K^{2}\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]$
+$=4|X_{0}|^{2}+4\left[T(1+M^{4})K^{2}+4M^{2}K^{2}\right]\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]$
+$\leq 4|X_{0}|^{2}+8\left[T(1+M^{4})K^{2}+4M^{2}K^{2}\right]\left(T+\int_{0}^{T}\mathbb{E}[|X_{s}^{\theta}|^{2}]ds\right).$
+By Gronwall’s inequality, there exists $C>0$ such that:
+$\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{\theta}|^{2}\right]\leq C(1+|X_{0}|^{2})e^{CT}.$
+Crucially, the constant $C$ depends only on $K$, $M$, and $T$, not on the specific control $\theta$.
+Step 5: Uniqueness. If $X_{t}^{\theta}$ and $Y_{t}^{\theta}$ are two solutions, then:
+$\mathbb{E}[|X_{t}^{\theta}-Y_{t}^{\theta}|^{2}]\leq 3(T+TM^{4}+M^{2})K^{2}\int_{0}^{t}\mathbb{E}[|X_{s}^{\theta}-Y_{s}^{\theta}|^{2}]ds.$
+By Gronwall’s inequality, $\mathbb{E}[|X_{t}^{\theta}-Y_{t}^{\theta}|^{2}]=0$ for all $t$. ∎
+Theorem 4.2 (G-SDE Solution via Parameterization).
 Under the same Lipschitz and growth conditions, the G-SDE (4.2) admits a unique solution given by:
@@
 The solution is constructed in three steps:
-  1. 1.
-    For each $\theta\in\mathcal{A}^{\Theta}$, solve the classical SDE (4.1)
-  2. 2.
-    The boundedness of $\Theta$ ensures uniform estimates:
+1. For each $\theta\in\mathcal{A}^{\Theta}$, solve the classical SDE (4.1)
+2. The boundedness of $\Theta$ ensures uniform estimates:
 $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{s}^{\theta}|^{2}\right]\leq C(1+|X_{0}|^{2})e^{Ct}$
-  3. 3.
-    The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation
+3. The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation ∎
 The compactness of $\Theta$ is crucial for preserving the solution properties:
@@
   • Maintaining moment estimates and path regularity
-This framework provides a rigorous bridge from classical SDE theory to G-SDEs, leveraging the bounded control structure to transfer solution properties.
 The G-SDE solution is reconstructed from the parameterized solutions:
@@
 The solution inherits properties from the parameterized solutions:
-Theorem 4.4 (Properties of Solutions ).
-The G-SDE solution satisfies:
-  1. 1.
-    Moment estimates: $\hat{\mathbb{E}}\left[\sup_{0\leq s\leq t}|X_{s}|^{p}\right]<\infty$ for $p\geq 2$
-  2. 2.
-    Continuous dependence: If $b_{n}\to b$, $\sigma_{n}\to\sigma$ uniformly, then $X_{t}^{n}\to X_{t}$ in $L_{G}^{2}$
-4.3 Geometric Brownian Motion — A Clear Example
+Remark 4.1.
+For comparison theorems of G-SDEs, we can directly draw on well-established results from classical SDE theory through our parametric representation.
+Similarly, a priori estimates and stability analysis for G-SDEs can be obtained by applying the corresponding classical results to each parameterized SDE in the family, provided we maintain uniform estimates with respect to the control parameters.
+This approach leverages the extensive existing literature on classical stochastic differential equations.
+Indeed, numerous fundamental properties and results for classical SDEs—including comparison theorems, a priori estimates, and stability analysis—can be systematically transferred to the G-SDE framework through our parametric representation.
+The established estimates in classical SDE theory demonstrate consistent behavior across control parameters, making them well-suited for application within our unified framework.
+4.2 Geometric Brownian Motion
 Consider the geometric-type G-SDE with volatility uncertainty:
 $dX_{t}=\sigma X_{t}dB_{t}+\gamma X_{t}d\langle B\rangle_{t},\quad X_{0}=x_{0}>0$ (4.7)
-Through the isomorphism $\Phi$, we transform the G-SDE into a family of classical SDEs:
+[…]
 $\Phi(X_{t})=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
@@
 $X_{t}^{\theta}=x_{0}\exp\left(\sigma\int_{0}^{t}\theta_{s}dW_{s}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)\int_{0}^{t}\theta_{s}^{2}ds\right)$ (4.14)
-Note that the two are identical under isomorphism.
+Note that the two are identical under homomorphism.
 The G-expectation is obtained by taking supremum over controls:
@@
 $\hat{\mathbb{E}}[X_{t}]=\sup_{\theta\in[\underline{\sigma},\overline{\sigma}]}x_{0}\exp(\gamma\theta^{2}t)=\begin{cases}x_{0}\exp(\gamma\overline{\sigma}^{2}t),&\gamma>0\\ x_{0}\exp(\gamma\underline{\sigma}^{2}t),&\gamma<0\end{cases}$ (4.16)
-5 Properties of G-Martingales
-In G-expectation theory, martingales exhibit fundamentally different behaviors due to volatility uncertainty.
-This chapter explains symmetric and asymmetric martingales through the lens of parameterized controls, revealing their economic interpretations as different types of ”fair gambles”.
-Under the isomorphism $\Phi$, a G-martingale $M_{t}$ is represented as a family of classical processes:
-$\Phi(M_{t})=\{M_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
-where each $M_{t}^{\theta}$ is adapted to the filtration generated by $W_{t}^{\theta}$.
-5.1 Symmetric Martingales: Fair Gambles in All Scenarios
-Definition 5.1 (Symmetric Martingale).
-A process $M_{t}$ is a symmetric G-martingale if for every control $\theta\in\mathcal{A}_{s,t}^{\Theta}$, $M_{t}^{\theta}$ is a classical martingale under $\mathbb{P}$:
-$\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta}|\mathcal{F}_{s}]=M_{s}^{\theta}\quad\forall s\leq t.$
-[…]
-5.2 Asymmetric Martingales: Favorable Gambles with Worst-Case Fairness
-Definition 5.2 (Asymmetric Martingale).
-A process $M_{t}$ is an asymmetric G-martingale if:
+5 Parametric Approach to G-Backed Stochastic Differential Equations
+5.1 Properties of G-Martingales
+In $G$-expectation theory, martingales are classified based on their behavior under different volatility scenarios.
+Through the parametric representation, we identify two fundamental types of $G$-martingales.
+Symmetric martingales represent processes that remain fair games regardless of the volatility environment.
+In every possible volatility scenario, these processes maintain their martingale property, making them truly robust fair gambles.
+Example: The $G$-Brownian motion $B_{t}$ is a symmetric martingale, as it behaves as a fair game under all volatility controls.
+Asymmetric martingales describe processes that are generally favorable—never expected to lose in any volatility scenario—but become exactly fair only in the worst-case volatility environment.
+These represent gambles with built-in protection against losses while maintaining fairness under the most adverse conditions.
+Example: The quadratic variation process $\langle B\rangle_{t}-\overline{\sigma}^{2}t$ is an asymmetric martingale, favorable in most scenarios but fair under maximum volatility.
+This classification reveals how volatility uncertainty creates distinct types of ”fair” processes, with symmetric martingales providing universal fairness and asymmetric martingales offering protection while maintaining worst-case fairness.
+5.2 The Emergence of $K_{t}$ A Cylindrical Perspective
+Consider a cylindrical Brownian motion with terminal condition $\xi=\phi(B_{t_{1}},B_{t_{2}}-B_{t_{1}})$ for $0<t_{1}<t_{2}\leq T$.
+Here, both generators f and g are taken to be identically zero.The global robust valuation at time 0 is achieved by an optimal control $\theta^{*}=(\theta^{*}_{1},\theta^{*}_{2})$ over the entire horizon $[0,T]$:
+$Y_{0}=\mathbb{E}_{\mathbb{P}}[\xi^{\theta^{*}}]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}[\xi^{\theta}]$
+At time $t_{1}$, the path $B_{t_{1}}$ is realized, and the control from 0 to $t_{1}$ is fixed at some value $\theta_{1}$.
+The conditional valuation then allows reoptimization over the remaining horizon $t_{1},T]$ with a new control $\theta_{2}$:
+$Y_{t_{1}}^{\theta_{1}}=\sup_{\theta_{2}\in\mathcal{A}_{t_{1},T}^{\Theta}}\mathbb{E}_{\mathbb{P}}[\xi^{\theta_{1},\theta_{2}}|\mathcal{F}_{t_{1}}]$
+The critical issue arises because the fixed control $\theta_{1}$ in the initial segment may not align with the global optimal control $\theta^{*}_{1}$.
+Even with optimal reoptimization over $\theta_{2}$ in the second segment, the local valuation $Y_{t_{1}}^{\theta_{1}}$ can never exceed the conditional expectation under the global optimum $\theta^{*}$.
+This discrepancy stems from the fact that the global optimal control coordinates both segments simultaneously, while the local valuation is constrained by the fixed and potentially suboptimal $\theta_{1}$.
+To bridge this gap, the compensation process $K_{t}$ is introduced.
+It quantifies the loss in valuation due to the suboptimal fixed control in the initial segment.
+Specifically, $K_{t}$ ensures temporal consistency by compensating for the difference between the local valuation under $\theta_{1}$ and the global optimal valuation.
+Importantly, $K_{t}$ vanishes only when the fixed control $\theta_{1}$ matches the global optimal $\theta^{*}_{1}$, indicating that the process behaves as a martingale solely under the global optimum.
+This vanishing property defines the asymmetric martingale structure, where the compensation persists under suboptimal controls but disappears under optimal coordination.
+5.3 From Classical Valuation to Robust Pricing under Uncertainty
+In classical stochastic analysis, the solution $Y_{t}$ of a Backward Stochastic Differential Equation (BSDE) represents the fair value of a contingent claim at time $t$.
+This valuation has a natural representation in terms of conditional expectations.
+5.3.1 Classical BSDE: Conditional Expectation Representation
+Consider the classical BSDE:
+$\begin{cases}-dY_{t}=f(t,Y_{t},Z_{t})dt-Z_{t}dW_{t}\\ Y_{T}=\xi\end{cases}$ (5.1)
+Under appropriate conditions, the solution admits the representation:
+$Y_{t}=\mathbb{E}\left[\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds\bigg|\mathcal{F}_{t}\right]$ (5.2)
+This representation reveals the fundamental economic interpretation: at any time $t$, the value $Y_{t}$ equals the conditional expectation of the terminal payoff $\xi$ plus the accumulated ”cost” or ”benefit” from the generator $f$ over the remaining period $[t,T]$.
+5.3.2 Transition to G-BSDE: Incorporating Model Uncertainty
+When we introduce model uncertainty, particularly volatility ambiguity, we transition to the framework of G-expectation. The G-BSDE takes the form:
+$\begin{cases}-dY_{t}=f(t,Y_{t},Z_{t})dt+g(t,Y_{t},Z_{t})d\langle B\rangle_{t}-Z_{t}dB_{t}-dK_{t}\\ Y_{T}=\xi\end{cases}$ (5.3)
+where $B_{t}$ is a G-Brownian motion, and $K_{t}$ is a continuous non-decreasing process with $K_{0}=0$.
+5.4 Conditional Expectation Forms and the Emergence of the Compensation Process
+5.4.1 G-BSDE as Robust Conditional Expectation
+The $G$-BSDE solution admits a representation in terms of the robust conditional expectation:
+$Y_{t}=\hat{\mathbb{E}}_{t}\left[\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds+\int_{t}^{T}g(s,Y_{s},Z_{s})d\langle B\rangle_{s}\right]$ (5.4)
+5.4.2 The Fundamental Distinction from Classical BSDE Parameterization
+A critical observation emerges when we contrast this robust formulation with a naive parameterization of classical BSDE solutions.
+If we were to simply consider a family of classical BSDEs parameterized by control processes $\theta\in\mathcal{A}^{\Theta}$:
+$Y_{t}^{\theta}=\mathbb{E}_{t}\left[\xi^{\theta}+\int_{t}^{T}f(s,Y_{s}^{\theta},Z_{s}^{\theta})ds+\int_{t}^{T}g(s,Y_{s}^{\theta},Z_{s}^{\theta})\theta_{s}^{2}ds\right]$
+we would encounter a fundamental temporal inconsistency.
+The valuation $Y_{t}^{\theta}$ for a fixed control $\theta$ inherently depends on the entire future trajectory of the control process $\{\theta_{s}\}_{s\in[t,T]}$.
+This violates the principle of robust pricing under uncertainty, where at time $t$, the agent cannot foresee which volatility scenario will materialize in the future.
+The genuine robust valuation $Y_{t}$ in the $G$-framework must satisfy a more stringent requirement: it represents the worst-case value without clairvoyance.
+At each time $t$, the agent evaluates the claim under the assumption that future volatility will evolve in the most adverse manner possible, but crucially, the agent cannot commit to a specific future control path since this would require knowledge of future uncertainty realizations.
+This distinction is precisely why the compensation process $K_{t}$ emerges inevitably in the $G$-BSDE framework.
+The process $K_{t}$ compensates for the inability to pre-commit to a single worst-case volatility path, capturing the economic cost of maintaining robustness against all possible future volatility scenarios while remaining adapted to the filtration.
+In essence, while a parameterized family of classical BSDE solutions $\{Y_{t}^{\theta}\}$ represents valuations under predetermined volatility paths, the $G$-BSDE solution $Y_{t}$ embodies a genuinely robust valuation that dynamically adjusts to the worst-case scenario without presuming knowledge of future volatility realizations.
+5.5 Parametric Representation of G-BSDEs
+Building upon the parametric framework established in previous sections, we extend this methodology to the analysis of $G$-BSDEs.
+In this approach, each component of the $G$-BSDE solution is represented as a family of processes parameterized by volatility controls:
+$Y_{t}$ $=\{Y_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}},\quad Z_{t}=\{Z_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}},\quad K_{t}=\{K_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
+For each fixed control process $\theta\in\mathcal{A}^{\Theta}$, the triple $(Y^{\theta},Z^{\theta},K^{\theta})$ satisfies an equation within the linear expectation framework:
+$Y_{t}^{\theta}=\xi^{\theta}+\int_{t}^{T}\left[f(s,Y_{s}^{\theta},Z_{s}^{\theta})+g(s,Y_{s}^{\theta},Z_{s}^{\theta})\theta_{s}^{2}\right]ds-\int_{t}^{T}Z_{s}^{\theta}\theta_{s}dW_{s}+(K_{T}^{\theta}-K_{t}^{\theta})$ (5.5)
+This representation enables us to leverage classical stochastic analysis techniques for conducting a priori estimates and stability analysis.
+The essential requirement is that all estimates must hold uniformly with respect to the control parameter $\theta\in\mathcal{A}^{\Theta}$.
+Following the established methodology in classical GBSDE theory, we begin with terminal conditions given by cylindrical Brownian motions.
+Through appropriate approximation schemes and convergence arguments, we extend the results to general terminal conditions.
+The uniform estimates with respect to the control parameter ensure the consistency of this approach and allow for the construction of $G$-BSDE solutions through limiting procedures.
+6 The Essence of Homomorphism: A Unified Perspective on Probabilistic Lenses
+In this section, we reveal the fundamental nature of the homomorphic representation by demonstrating that we are always working with the same stochastic variable, merely viewing it through different probabilistic lenses on the canonical path space $C([0,1])$.
+6.1 The Probabilistic Perspective of the Homomorphism
+We begin with the canonical path space $\Omega=C([0,1])$ endowed with the topology of uniform convergence.
+The fundamental insight is that we are working with a single, unified stochastic framework where the canonical process $B_{t}(\omega)=\omega(t)$ remains the same mathematical object throughout our analysis.
+The apparent multiplicity arises from the fact that this same canonical process can be interpreted through different probabilistic lenses.
+Each admissible control $\theta\in\mathcal{A}^{\Theta}$ provides a different perspective on the same underlying reality:
+$B_{t}=\int_{0}^{t}\theta_{s}dW_{s}^{\theta}\quad\text{under }P^{\theta},$
+where $W^{\theta}$ is a standard Brownian motion under $P^{\theta}$.
+In this framework, the canonical process serves as the paradigmatic representative of stochastic integrals across all probabilistic perspectives.
+Crucially, we are not mapping between different stochastic variables, but rather interpreting the same canonical process through different volatility characteristics.
+The homomorphism $\Phi$ is not a mapping between different mathematical objects, but rather a device that organizes our different perspectives on the same stochastic reality.
+For a cylinder function $\xi=\varphi(B_{t_{1}},\dots,B_{t_{n}})$, we have:
+$\Phi(\xi)=\left\{\xi^{\theta}=\varphi(B_{t_{1}},\dots,B_{t_{n}})\right\}_{\theta\in\mathcal{A}^{\Theta}}.$
+The profound insight is that $\xi^{\theta}$ is not a different random variable from $\xi$, but rather the same random variable $\xi$ viewed through the probabilistic lens of $P^{\theta}$.
+The parameterized family $\{\xi^{\theta}\}$ represents the different interpretations of the same underlying cylinder function across various volatility scenarios.
+The homomorphism $\Phi$ reveals its true nature not as a mapping between different mathematical objects, but as an organizational framework that:
+1. Unifies Perspectives: It demonstrates how the same stochastic object can be consistently interpreted through different probabilistic lenses.
+2. Maintains Identity: Through uniform convergence, it ensures that the stochastic object maintains its identity across all interpretations.
+3. Enables Robust Analysis: It provides a foundation for conducting stochastic analysis that is robust to model uncertainty, as we are always working with the same mathematical objects.
+The parameterized family $\{X^{\theta}\}$ is not a collection of different random variables, but rather a representation of the different interpretations of the same stochastic variable $X$ across various volatility scenarios.
+We have elucidated the essence of the homomorphic representation as a framework for understanding how the same stochastic object can be viewed through different probabilistic lenses on the canonical path space.
+The key insight is that we are always working with the same mathematical objects; the homomorphism simply organizes the different ways of interpreting these objects under model uncertainty.
+This approach provides not only a powerful analytical tool but also a profound philosophical perspective on the nature of stochastic objects: they are entities that maintain their identity regardless of the probabilistic lens through which we view them.
+6.2 Uniform Convergence and Mathematical Coherence
+When extending from cylinder functions to general $G$-random variables, we employ approximation techniques.
+Let $X\in L^{p}_{G}(\Omega)$ and let $\{\xi_{n}\}$ be a sequence of cylinder functions converging to $X$ in $L^{p}_{G}$.
+By the isometry property, we have:
+$\sup_{\theta\in\mathcal{A}^{\Theta}}\|X-\xi_{n}\|_{L^{p}(P^{\theta})}\to 0.$
+The uniformity of this convergence is essential for maintaining the coherence of our framework.
+Without it, the interpretation of the limiting object $X$ could become fragmented across different probabilistic perspectives, with each $\theta$ requiring a different approximating subsequence.
+However, the uniform convergence ensures that we can select a single subsequence $\{\xi_{n_{k}}\}$ that works simultaneously for all $\theta\in\mathcal{A}^{\Theta}$:
+$\xi_{n_{k}}\to X\quad\text{quasi-surely for all }\theta\in\mathcal{A}^{\Theta}.$
+This simultaneous convergence guarantees that $X$ maintains its identity as a single, coherent stochastic variable across all probabilistic perspectives.
+The different interpretations $X$ under different $P^{\theta}$ are not different mathematical objects, but rather different perspectives on the same underlying reality.
+6.3 Quasi-Sure Inheritance of Path Properties
+In our parametric framework, if for each $\theta\in\mathcal{A}^{\Theta}$ the process $X^{\theta}$ satisfies a path property almost surely under $P^{\theta}$, then $X$ satisfies this property $\mathcal{P}$-quasi-surely.
+For example, if each $X^{\theta}$ is Hölder continuous with exponent $\alpha$ $P^{\theta}$-almost surely, then $X$ is Hölder continuous with exponent $\alpha$ $\mathcal{P}$-quasi-surely.
+This direct inheritance applies to all path properties. By verifying properties individually under each probabilistic perspective, we automatically obtain robust quasi-sure results.
+Remark 6.1.
+It is important to distinguish between the path properties discussed here and the quasi-sure continuity in classical $G$-expectation theory.
+The latter typically refers to continuity with respect to the topology of $C([0,1])$ space under capacity.
+Here we focus on intrinsic path properties such as Hölder continuity, boundedness, and differentiability.
+We note that within the $L^{p}_{G}$ space framework, the quasi-sure continuity property of the $G$-framework is naturally satisfied.
+These are distinct concepts: quasi-sure continuity concerns the topological structure of the path space, while the path properties mentioned here are analytical characteristics of individual sample paths.
+6.4 The Constructive Philosophy of GGExpectation: From Canonical Space to Robust Representatives
+The construction of $G$-expectation theory establishes a multiple probability measures framework on the canonical continuous function space $C[0,1]$.
+The following diagram illustrates this constructive process:
+Canonical Path Space $C[0,1]$ with coordinate process $B_{t}(\omega)=\omega(t)$ Deterministic mathematical object Attach Probability Measure Family $\{P_{\theta}\}_{\theta\in\Theta}$ Each $\theta$ provides a distinct probabilistic perspective Cylinder Functionals Natural definition: $\xi=\varphi(B_{t_{1}},\ldots,B_{t_{n}})$ Well-defined under all $P_{\theta}$ simultaneously $L_{G}^{p}$ Completion Uniform convergence: $\sup_{\theta}\|X-\xi_{n}\|_{L^{p}(P_{\theta})}\to 0$ Selects robust representatives across all measures Robust Analysis Framework Quasi-sure properties preserved Divide-and-conquer philosophy enabled Key Insight 1 Same mathematical object $B_{t}$ Different probabilistic interpretations $B_{t}=\int_{0}^{t}\theta_{s}dW_{s}^{\theta}$ under $P_{\theta}$ Key Insight 2 $L_{G}^{p}$ ensures same subsequence Works for all $P_{\theta}$ simultaneously Solves representation ambiguity problem Key Insight 3 Analyze under each $P_{\theta}$ Integrate via uniform estimates True divide-and-conquer approach
+6.4.1 Detailed Construction Process
+Stage 1: Foundation – Canonical Space with Multiple Perspectives
+We begin with the canonical path space $(\Omega,\mathcal{F})=(C[0,1],\mathcal{B}(C[0,1]))$, where the coordinate process $B_{t}(\omega)=\omega(t)$ is a purely mathematical object.
+For each sample path $\omega\in C[0,1]$, the trajectory $B_{t}(\omega)$ is completely determined and represents a continuous function on $[0,1]$.
+The crucial innovation is attaching a family of probability measures $\{P_{\theta}\}_{\theta\in\Theta}$ to this space.
+Each control process $\theta\in\mathcal{A}^{\Theta}$ induces a probability measure $P_{\theta}$ under which the coordinate process satisfies:
+$B_{t}=\int_{0}^{t}\theta_{s}dW_{s}^{\theta}\quad\text{under }P_{\theta}$
+where $W_{t}^{\theta}$ is a standard Brownian motion under $P_{\theta}$.
+This means the same mathematical entity $B_{t}$ acquires different probabilistic interpretations under different measures $P_{\theta}$.
+Stage 2: Natural Objects – Cylinder Functionals
+For cylinder functionals of the form $\xi=\varphi(B_{t_{1}},\ldots,B_{t_{n}})$, the definition is intrinsically natural:
+$\xi^{\theta}=\varphi(B_{t_{1}},\ldots,B_{t_{n}})\quad\text{under each }P_{\theta}$
+Since cylinder functionals depend only on finitely many time points, they are well-defined independently of any specific probability measure. This makes them ideal building blocks for our theory.
+Stage 3: Completion – Resolving Representation Ambiguity
+When we complete the space of cylinder functionals using the $L_{G}^{p}$ norm, we encounter a fundamental challenge: in classical probability theory, $L^{p}$ limits are defined modulo null sets, and different probability measures may have different null sets.
+The $L_{G}^{p}$ norm brilliantly resolves this through the uniform convergence condition:
+$\sup_{\theta\in\mathcal{A}^{\Theta}}\|X-\xi_{n}\|_{L^{p}(P_{\theta})}\to 0$
+This ensures we can select a single subsequence $\{\xi_{n_{k}}\}$ that converges appropriately under all measures $P_{\theta}$ simultaneously.
+The limit thus defines a robust representative that behaves consistently across all probabilistic perspectives.
+Stage 4: Analysis – The Divide-and-Conquer Philosophy
+The completed framework enables powerful analytical strategies:
+- Quasi-sure properties: Path properties that hold $P_{\theta}$-almost surely for all $\theta$ become quasi-sure properties in the $G$-framework.
+- Probabilistic lens analysis: We can safely analyze our objects under each fixed $P_{\theta}$, leveraging classical stochastic calculus tools.
+- Uniform integration: Results from different perspectives are integrated through supremum operations and uniform estimates.
+6.5 The Ontology of BSDE Solutions Under Model Uncertainty: Intrinsic Identification via Quadratic Variation
+Building upon our analysis of the canonical space and probabilistic lenses, we now examine the fundamental structure of BSDE solutions under model uncertainty through an ontological perspective.
+This inquiry addresses not merely technical aspects, but the very nature of existence for stochastic processes in the presence of uncertainty.
+Consider the standard BSDE under model uncertainty:
+$Y_{t}=\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})ds+\int_{t}^{T}g(s,Y_{s},Z_{s})d\langle B\rangle_{s}-\int_{t}^{T}Z_{s}dB_{s}-(K_{T}-K_{t})$ (6.1)
+Within our parametric framework, the terminal condition $\xi$ exists as an objective entity on the canonical path space.
+Through the dynamic programming principle, the value process $Y_{t}$ similarly emerges as an independently existing mathematical object, defined in a robust sense on the canonical space.
+However, the ontological status of the diffusion term $Z_{t}$ and compensation term $K_{t}$ appears less immediately evident.
+Through our homomorphic mapping, they initially manifest as parameterized families:
+$\Phi(Z)\sim\{Z^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}},\quad\Phi(K)\sim\{K^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ (6.2)
+This perspective raises a fundamental question: can these families be aggregated into a single, objectively existing mathematical entity?
+6.5.1 Quadratic Variation as an Anchor of Objectivity
+The quadratic variation process $\langle B\rangle_{t}$ stands as one of the most rigid structures in stochastic analysis.
+In our framework, it manifests as an intrinsic, robust process defined through pathwise limits, independent of any specific probability measure.
+If we accept $Y_{t}$ as an objectively existing semimartingale, then its quadratic variation $\langle Y\rangle_{t}$ must similarly exist as an objective mathematical entity.
+Examining the dynamics of the BSDE:
+$dY_{t}=\text{(finite variation drift terms)}dt+Z_{t}dB_{t}+dK_{t}$ (6.3)
+We observe that $K_{t}$, being a non-decreasing compensation process, contributes only finite variation.
+The classical drift terms and terms involving $d\langle B\rangle$ similarly contribute finite variation.
+Consequently, the entire quadratic variation of $Y$ must originate exclusively from the stochastic integral term $Z_{t}dB_{t}$.
+Under any fixed probabilistic perspective $P^{\theta}$, we have the classical relationship:
+$d\langle Y^{\theta}\rangle_{t}=(Z_{t}^{\theta})^{2}d\langle B^{\theta}\rangle_{t}=(Z_{t}^{\theta})^{2}\theta_{t}^{2}dt$ (6.4)
+This reveals that, across all perspectives, the square of $Z$ functions as the Radon-Nikodym derivative relating the quadratic variations:
+$(Z_{t}^{\theta})^{2}=\frac{d\langle Y^{\theta}\rangle_{t}}{d\langle B^{\theta}\rangle_{t}}$ (6.5)
+6.5.2 From Multiple Perspectives to Unified Entity
+This relationship yields a profound conclusion: if $Y$ and $B$ are objective entities, then their quadratic variations $\langle Y\rangle$ and $\langle B\rangle$ must also be objective.
+Therefore, the process $Z_{t}$, defined as:
+$Z_{t}:=\sqrt{\frac{d\langle Y\rangle_{t}}{d\langle B\rangle_{t}}}$ (6.6)
+must necessarily exist as a unified, objective mathematical entity.
+The parameterized family $\{Z^{\theta}\}$ is not arbitrary but subject to an intrinsic constraint: its members must be different manifestations of the same fundamental object under different probabilistic lenses.
+With the objective status of $Y_{t}$ and $Z_{t}$ established, the compensation process $K_{t}$ follows necessarily. Rearranging the BSDE:
+$K_{T}-K_{t}=\xi-Y_{t}+\int_{t}^{T}f(s,Y_{s},Z_{s})ds+\int_{t}^{T}g(s,Y_{s},Z_{s})d\langle B\rangle_{s}-\int_{t}^{T}Z_{s}dB_{s}$ (6.7)
+6.6 Geometric Necessity: From Classical Foundations to Intrinsic Objectivity
+The insight provided by quadratic variation leads us to a powerful conclusion concerning the nature of BSDE solutions under model uncertainty: their constituents are bound not just by probabilistic coherence, but by rigid geometric constraints on the canonical path space.
+The construction begins with the parameterized family $\{Y_{t}^{\theta}\}$, obtained by solving classical BSDEs under each fixed scenario $\theta$.
+Establishing strict uniform convergence for this family over the parameter space $\Theta$ is essential, as it ensures the stability and robustness of the limiting process.
+However, convergence alone does not automatically yield the final object.
+It is through the application of the foundational methodology from the seminal work of El Karoui, Peng, and Quenez [41] that we synthesize these convergent paths to form the genuine process $Y_{t}$.
+This technique allows us to construct $Y_{t}$ not merely as a limit point, but as a unified, objectively existing mathematical entity on the canonical space, defined in a robust sense.
+Once the objective existence of $Y_{t}$ is established through this synthesis of uniform convergence and classical construction techniques, the determination of the remaining components $(Z_{t},K_{t})$ transitions from a probabilistic selection problem to a matter of geometric necessity.
+The fundamental relationship,
+$Z_{t}^{2}=\frac{d\langle Y\rangle_{t}}{d\langle B\rangle_{t}},$
+imposes a geometric constraint.
+It dictates that the infinitesimal ”roughness” of the path $Y_{t}$ must be entirely derivable from the roughness of the driving Brownian motion $B_{t}$, scaled locally by $Z_{t}^{2}$.
+The finite variation components are geometrically ”smooth” and cannot contribute to this roughness.
+Consequently, the apparent multiplicity of perspectives for the diffusion term is collapsed by this geometric rigidity.
+The parameterized family $\{Z^{\theta}\}$ cannot be arbitrary but must conform to this intrinsic ratio, forcing its coalescence into the single, objective entity $Z_{t}$.
+With $Y_{t}$ constructed via the methodology of [41] and $Z_{t}$ determined by geometric necessity, the compensation process $K_{t}$ emerges as the unique residual required to balance the dynamical equation.
+The entire solution triplet $(Y,Z,K)$ thus constitutes an objectively existing entity, bound by the indissoluble geometric structure of the canonical path space.
+Remark 6.2.
+The geometric constraint perspective introduced here provides a foundational principle for the robust analysis of G-BSDEs.
+This approach reveals a crucial methodological insight: once the primary value process $Y$ is established as a robust object through uniform convergence and classical construction techniques, the remaining components $Z$ and $K$ are automatically determined through geometric necessity.
+This means that ensuring the robustness of $Y$ automatically guarantees the robustness of the entire solution triplet $(Y,Z,K)$, providing a powerful unifying principle for the analysis of backward stochastic systems under model uncertainty.
+7 Comparative Analysis of Estimation Precision
+In stochastic analysis under model uncertainty, the practical utility of theoretical frameworks depends critically on the precision of their estimates.
+A significant challenge within the traditional $G$-expectation framework involves the tendency of its fundamental estimates to be overly conservative.
+This conservatism arises not merely from computational approximations but from the inherent structure of the ”stepwise optimization” paradigm employed in establishing foundational inequalities.
+In this section, we demonstrate that the parametric representation framework addresses this structural limitation.
+By shifting the analytical approach, our methodology moves from providing loose upper bounds to delivering sharp, and in fundamental cases, exact estimates.
+We illustrate this through a rigorous comparison of estimations for the second moment of the quadratic variation of $G$-Brownian motion.
+7.1 he Conservative Upper-Bound Bias in the G-Framework: A Methodological Renovation
+We investigate the estimation of $\hat{\mathbb{E}}[\langle B\rangle_{t}^{2}]$, a fundamental quantity in the analysis of the variance of volatility.
+We first present the standard bound derived within the original $G$-expectation framework, followed by the exact calculation achievable through our parametric approach.
+7.1.1 Estimation via the Traditional Framework
+The traditional approach relies on decomposing the target random variable and applying sublinear expectations step-by-step.
+This process often necessitates the use of inequalities and heavily breaches the correlation structure between variables due to the subadditivity property: $\hat{\mathbb{E}}[X+Y]\leq\hat{\mathbb{E}}[X]+\hat{\mathbb{E}}[Y]$.
+The following lemma restates a known bound derived using standard techniques in $G$-expectation theory [1]
+Lemma 7.1 (Traditional Upper Bound).
+Let $\langle B\rangle_{t}$ be the quadratic variation process of a $G$-Brownian motion $B_{t}$ with volatility uncertainty interval $\underline{\sigma},\bar{\sigma}]$. The following upper bound holds:
+$\hat{\mathbb{E}}\left[\langle B\rangle_{t}^{2}\right]\leq 10\bar{\sigma}^{4}t^{2}.$ (7.1)
+Proof.
+We utilize the identity $\langle B\rangle_{t}=B_{t}^{2}-2\int_{0}^{t}B_{u}dB_{u}$.
+Applying the sublinearity of $\hat{\mathbb{E}}$ and the elementary inequality $(a-b)^{2}\leq 2a^{2}+2b^{2}$, we obtain:
+$\hat{\mathbb{E}}\left[\langle B\rangle_{t}^{2}\right]$ $=\hat{\mathbb{E}}\left[\left(B_{t}^{2}-2\int_{0}^{t}B_{u}dB_{u}\right)^{2}\right]$
+$\leq\hat{\mathbb{E}}\left[2B_{t}^{4}+8\left(\int_{0}^{t}B_{u}dB_{u}\right)^{2}\right].$
+A critical step involves applying subadditivity, which inherently decouples the two terms, implicitly assuming a worst-case scenario for each term independently:
+$\hat{\mathbb{E}}\left[\langle B\rangle_{t}^{2}\right]\leq 2\hat{\mathbb{E}}\left[B_{t}^{4}\right]+8\hat{\mathbb{E}}\left[\left(\int_{0}^{t}B_{u}dB_{u}\right)^{2}\right].$ (7.2)
+Standard results for $G$-Brownian motion give $\hat{\mathbb{E}}[B_{t}^{4}]=3\bar{\sigma}^{4}t^{2}$. For the second term, applying the inequality for $G$-expectations yields:
+$8\hat{\mathbb{E}}\left[\left(\int_{0}^{t}B_{u}dB_{u}\right)^{2}\right]$ $\leq 8\bar{\sigma}^{2}\hat{\mathbb{E}}\left[\int_{0}^{t}B_{u}^{2}du\right]$
+$\leq 8\bar{\sigma}^{2}\int_{0}^{t}\hat{\mathbb{E}}\left[B_{u}^{2}\right]du$
+$=8\bar{\sigma}^{2}\int_{0}^{t}\bar{\sigma}^{2}udu=4\bar{\sigma}^{4}t^{2}.$
+Substituting these back into (7.2):
+$\hat{\mathbb{E}}\left[\langle B\rangle_{t}^{2}\right]\leq 2(3\bar{\sigma}^{4}t^{2})+4\bar{\sigma}^{4}t^{2}=10\bar{\sigma}^{4}t^{2}.$
+This concludes the proof of the bound. ∎
+7.1.2 Exact Calculation via Parametric Representation
+The parametric framework allows us to compute expectations within the classical linear probability space before applying the supremum. This preserves the algebraic structure of the random variables.
+Theorem 7.2 (Exact Estimation).
+Under the same conditions as Lemma 7.1, the parametric representation methodology yields the exact value:
+$\hat{\mathbb{E}}\left[\langle B\rangle_{t}^{2}\right]=\bar{\sigma}^{4}t^{2}.$ (7.3)
+Proof.
+By the definition of the parametric homomorphism, the $G$-expectation is the supremum of classical expectations over the set of admissible controls $\mathcal{A}^{\Theta}$:
+$\hat{\mathbb{E}}\left[\langle B\rangle_{t}^{2}\right]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\left(\left\langle B^{\theta}\right\rangle_{t}\right)^{2}\right].$
+Under a fixed probability measure $\mathbb{P}$ and control $\theta\in\mathcal{A}^{\Theta}$, the quadratic variation process is given by $\left\langle B^{\theta}\right\rangle_{t}=\int_{0}^{t}\theta_{s}^{2}ds$.
+It is crucial to note that this is a pathwise defined random variable under $\mathbb{P}$.
+Since the control set is $\Theta=[\underline{\sigma},\bar{\sigma}]$, we have $\underline{\sigma}^{2}\leq\theta_{s}^{2}\leq\bar{\sigma}^{2}$ for all $s\in[0,t]$, $\mathbb{P}$-a.s.
+Therefore, we have the pathwise comparison:
+$\left(\int_{0}^{t}\theta_{s}^{2}ds\right)^{2}\leq\left(\int_{0}^{t}\bar{\sigma}^{2}ds\right)^{2}=(\bar{\sigma}^{2}t)^{2}=\bar{\sigma}^{4}t^{2},\quad\mathbb{P}\text{-a.s.}$
+Taking the expectation $\mathbb{E}_{\mathbb{P}}$ on both sides yields:
+$\mathbb{E}_{\mathbb{P}}\left[\left(\left\langle B^{\theta}\right\rangle_{t}\right)^{2}\right]\leq\bar{\sigma}^{4}t^{2}.$
+Since this upper bound is independent of $\theta$, it holds for the supremum. Furthermore, this bound is attained by the constant admissible control process $\theta_{s}\equiv\bar{\sigma}$. Thus:
+$\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\left(\int_{0}^{t}\theta_{s}^{2}ds\right)^{2}\right]=\bar{\sigma}^{4}t^{2}.$
+This proves that $\bar{\sigma}^{4}t^{2}$ is the statistically exact value under the worst-case scenario, not merely an upper bound. ∎
+Table 1 summarizes the discrepancy between the two approaches.
+Table 1: Comparison of Estimates for $\hat{\mathbb{E}}[\langle B\rangle_{t}^{2}]$
+Methodology Result Nature of Estimate
+Traditional $G$-Framework $10\bar{\sigma}^{4}t^{2}$ Conservative Upper Bound
+Parametric Representation $\bar{\sigma}^{4}t^{2}$ Exact Value
+The comparison reveals a tenfold difference in precision for this fundamental quantity.
+The traditional method provides a bound that is an order of magnitude larger than the true worst-case expectation realized by the optimal parameter path.
+7.2 Methodological Origins of Precision Differences
+The significant discrepancy in precision highlighted above is not accidental but stems from fundamentally different approaches to handling model uncertainty.
+The traditional framework operates under a paradigm that may be characterized as ”Pessimize step-by-step, then Calculate.” To ensure robustness, sublinear expectations are often applied sequentially to components of a complex random variable.
+For instance, when estimating $\hat{\mathbb{E}}[X+Y]$, the use of subadditivity $\hat{\mathbb{E}}[X]+\hat{\mathbb{E}}[Y]$ implicitly assumes the existence of a volatility scenario that is worst-case for $X$ and simultaneously worst-case for $Y$.
+In reality, the single market volatility path cannot simultaneously maximize two potentially conflicting functional forms.
+By breaking the dependence structure between $X$ and $Y$ through inequalities, this approach constructs a ”virtual” worst-case scenario that may exceed any physically realizable outcome.
+Conversely, the parametric approach follows a ”Calculate, then Pessimize globally” paradigm.
+For each fixed control process $\theta$, all calculations are performed within a classical linear probability space.
+In this phase, precise equalities (such as Itô isometries) hold, and the joint distribution and correlation structure of all variables are perfectly preserved.
+The element of uncertainty—the ”pessimization”—is introduced only in the final step via a single supremum operation over the space of admissible controls.
+This ensures that the final estimate corresponds to a coherent, physically realizable worst-case scenario.
+By avoiding intermediate inequalities and maintaining the integrity of the stochastic structure until the final aggregation step, the parametric methodology inherently minimizes estimation bias.
+7.3 Implications for Complex Stochastic Modeling
+While demonstrated here on a fundamental random variable, the precision gap has implications for more complex applications of $G$-expectation theory, particularly in areas such as quantitative finance and risk management.
+Financial models often involve complex, path-dependent functionals of the underlying processes (e.g., in exotic option pricing) or require multi-step backward induction (e.g., in dynamic risk measures or numerical schemes for G-BSDEs).
+In such contexts, the systemic bias introduced by stepwise conservative estimation can accumulate cumulatively or even geometrically across time steps or model components.
+If basic building blocks like quadratic variation estimates carry significant upward bias, models built upon them may yield valuations or risk capital requirements that are excessively conservative.
+The parametric representation offers a methodological pathway to mitigate this accumulated error.
+By ensuring that estimates remain exact or as sharp as possible at the fundamental level, it provides a basis for constructing robust models that reflect a more accurate assessment of worst-case model uncertainty, rather than artifacts of the estimation procedure itself.
+This suggests that rethinking the computational approach via parametric representations is crucial for deploying $G$-expectation theory in practical domains where precision is paramount.
+8 Unification of Theory and Extensions
+The parametric representation framework establishes a dedicated and rigorous connection between nonlinear expectation theory and classical stochastic analysis.
+This connection offers two primary mathematical advantages:
+1.
+Unification and Simplification: It allows for the systematic importation of established classical results into the $G$-framework, thereby unifying and simplifying the theoretical edifice.
+2.
+Expansion of Scope: It provides a viable pathway to integrate advanced analytical machinery, such as Malliavin calculus, with model uncertainty, opening frontiers of fine analysis that are inaccessible to the traditional paradigm.
+This chapter elucidates how the parametric methodology realizes these advantages, first by simplifying the proofs of fundamental inequalities and second by outlining future research directions in fine analysis.
+8.1 Unification via Linear Decomposition
+The core analytical principle of the parametric framework is the decomposition of nonlinear problems into linear execution phases followed by a uniform synthesis.
+This approach separates the analytical complexity into two distinct stages:
   1. 1.
-    For all $\theta\in\mathcal{A}_{s,t}^{\Theta}$, $M_{t}^{\theta}$ is a supermartingale under $\mathbb{P}$:
-    $\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta}|\mathcal{F}_{s}]\leq M_{s}^{\theta}\quad\forall s\leq t$
+    Linear Execution Phase: Stochastic analysis operations—such as proving inequalities, solving SDEs, and establishing estimates—are performed within the familiar, linear environment of classical probability theory for each fixed control parameter $\theta\in\mathcal{A}^{\Theta}$.
   2. 2.
-    There exists a worst-case scenario $\theta^{*}\in\Theta$ such that:
-    $\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta^{*}}|\mathcal{F}_{s}]=M_{s}^{\theta^{*}}$
-5.3 Key Properties
-[…]
+    Uniform Synthesis Phase: The results from the parameterized family are synthesized into the robust $G$-framework through the single operation of taking the supremum over the parameter space $\Theta$, provided uniform estimates are established.
+8.1.1 The Unification of Fundamental Inequalities
+We demonstrate the utility of this decomposition principle by unifying and simplifying the proofs of fundamental inequalities within the $G$-framework.
+The general procedure consists of three steps: Translation, Linear Execution, and Uniform Synthesis.
+Hölder Inequality Parametric Derivation:
+  1. 1.
+    Translation: For $X,Y\in L_{G}^{p}$, by definition of the homomorphism: $\hat{\mathbb{E}}[|XY|]=\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}Y^{\theta}\right|\right]$.
+  2. 2.
+    Linear Execution: For each fixed $\theta$, apply the classical Hölder inequality: $\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}Y^{\theta}\right|\right]\leq\left(\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}\right|^{p}\right]\right)^{1/p}\left(\mathbb{E}_{\mathbb{P}}\left[\left|Y^{\theta}\right|^{q}\right]\right)^{1/q}$, where $1/p+1/q=1$.
+  3. 3.
+    Uniform Synthesis: Taking the supremum over $\theta$ on both sides.
+Using the property $\sup_{\theta}(f(\theta)g(\theta))\leq\sup_{\theta}f(\theta)\sup_{\theta}g(\theta)$ for non-negative functions, we obtain the $G$-Hölder inequality:
+    $\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}Y^{\theta}\right|\right]\leq\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}\right|^{p}\right]\right)^{1/p}\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|Y^{\theta}\right|^{q}\right]\right)^{1/q}=\|\Phi(X)\|_{\mathcal{P}_{p}}\|\Phi(Y)\|_{\mathcal{P}_{q}}$
+Minkowski Inequality Parametric Derivation:
+  1. 1.
+    Translation: $\left(\hat{\mathbb{E}}\left[|X+Y|^{p}\right]\right)^{1/p}=\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}+Y^{\theta}\right|^{p}\right]\right)^{1/p}$.
+  2. 2.
+    Linear Execution: For each fixed $\theta$, apply the classical Minkowski inequality: $\left(\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}+Y^{\theta}\right|^{p}\right]\right)^{1/p}\leq\left(\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}\right|^{p}\right]\right)^{1/p}+\left(\mathbb{E}_{\mathbb{P}}\left[\left|Y^{\theta}\right|^{p}\right]\right)^{1/p}$.
+  3. 3.
+    Uniform Synthesis: Taking the supremum over $\theta$ preserves the inequality due to the subadditivity of the supremum operation:
+    $\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}+Y^{\theta}\right|^{p}\right]\right)^{1/p}\leq\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|X^{\theta}\right|^{p}\right]\right)^{1/p}+\left(\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|Y^{\theta}\right|^{p}\right]\right)^{1/p}$
+    This yields the $G$-Minkowski inequality: $\|X+Y\|_{L_{G}^{p}}\leq\|X\|_{L_{G}^{p}}+\|Y\|_{L_{G}^{p}}$.
+Burkholder-Davis-Gundy Inequality for $G$-Stochastic Integrals
+Classical Statement: For a $G$-Brownian motion $B_{t}$ and an adapted process $H_{t}$, the stochastic integral $\int_{0}^{t}H_{s}dB_{s}$ satisfies for $p>0$:
+$c_{p}\hat{\mathbb{E}}\left[\left(\int_{0}^{t}\left|H_{s}\right|^{2}d\langle B\rangle_{s}\right)^{p/2}\right]\leq\hat{\mathbb{E}}\left[\left|\int_{0}^{t}H_{s}dB_{s}\right|^{p}\right]\leq C_{p}\hat{\mathbb{E}}\left[\left(\int_{0}^{t}\left|H_{s}\right|^{2}d\langle B\rangle_{s}\right)^{p/2}\right]$
+where $c_{p},C_{p}>0$ are generic constants depending only on $p$.
+Parametric Derivation:
+  1. 1.
+    Translation: By the homomorphism $\Phi$ regarding Itô integrals and quadratic variations:
+    $\hat{\mathbb{E}}\left[\left|\int_{0}^{t}H_{s}dB_{s}\right|^{p}\right]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\left|\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right|^{p}\right]$
+    and
+    $\hat{\mathbb{E}}\left[\left(\int_{0}^{t}\left|H_{s}\right|^{2}d\langle B\rangle_{s}\right)^{p/2}\right]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\left(\int_{0}^{t}\left|H_{s}^{\theta}\right|^{2}\theta_{s}^{2}ds\right)^{p/2}\right]$
+  2. 2.
+    Linear Execution: For each fixed $\theta$, the process $M_{t}^{\theta}=\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}$ is a classical martingale with quadratic variation $\langle M^{\theta}\rangle_{t}=\int_{0}^{t}\left|H_{s}^{\theta}\theta_{s}\right|^{2}ds=\int_{0}^{t}\left|H_{s}^{\theta}\right|^{2}\theta_{s}^{2}ds$.
+The classical BDG inequality applies:
+    $c_{p}\mathbb{E}_{\mathbb{P}}\left[\langle M^{\theta}\rangle_{t}^{p/2}\right]\leq\mathbb{E}_{\mathbb{P}}\left[|M_{t}^{\theta}|^{p}\right]\leq C_{p}\mathbb{E}_{\mathbb{P}}\left[\langle M^{\theta}\rangle_{t}^{p/2}\right]$
+  3. 3.
+    Uniform Synthesis: Crucially, the constants $c_{p}$ and $C_{p}$ in the classical BDG inequality depend only on $p$ and are independent of the underlying probability measure or the specific filtration, thus they are independent of $\theta$.
+Taking the supremum over $\theta\in\mathcal{A}^{\Theta}$ on all sides yields the result directly.
+For the upper bound:
+    $\hat{\mathbb{E}}\left[\left|\int_{0}^{t}H_{s}dB_{s}\right|^{p}\right]$ $=\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left|M_{t}^{\theta}\right|^{p}\right]$
+    $\leq\sup_{\theta}\left(C_{p}\mathbb{E}_{\mathbb{P}}\left[\left(\int_{0}^{t}\left|H_{s}^{\theta}\right|^{2}\theta_{s}^{2}ds\right)^{p/2}\right]\right)$
+    $=C_{p}\sup_{\theta}\mathbb{E}_{\mathbb{P}}\left[\left(\int_{0}^{t}\left|H_{s}^{\theta}\right|^{2}\theta_{s}^{2}ds\right)^{p/2}\right]$
+    $=C_{p}\hat{\mathbb{E}}\left[\left(\int_{0}^{t}\left|H_{s}\right|^{2}d\langle B\rangle_{s}\right)^{p/2}\right]$
+    The lower bound follows identically.
+8.2 Extensions to Fine Analysis and Future Directions
+The parametric framework establishes a natural pathway to incorporate Malliavin calculus techniques into the analysis of $G$-expectations.
+This integration enables a refined sensitivity analysis by working parameter-wise within the Malliavin framework.
+By examining the Malliavin derivatives and other fine properties for each parameterized classical process $X^{\theta}$, our approach provides a systematic methodology for conducting robust sensitivity analysis under the $G$-expectation framework.
+This direction opens promising avenues for studying the regularity and sensitivity of solutions to stochastic systems in the presence of volatility uncertainty, which are currently difficult to access using traditional sublinear approaches.
 9 Conclusion
-This paper has established a fundamental and rigorous connection between the abstract theory of $G$-expectation and the concrete framework of parameterized stochastic control.
-By introducing the concept of controlled processes $\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$, we have constructed a complete isomorphism $\Phi$ that maps the nonlinear $G$-expectation space onto a family of classical probability spaces indexed by volatility controls.
-The core contribution of this work is threefold:
-  1. 1.
-    Unified Representation: We demonstrated that $G$-Brownian motion and $G$-SDE solutions admit a natural representation as suprema over parameterized classical processes, providing an intuitive and computationally tractable interpretation of model uncertainty.
-  2. 2.
-    Algebraic and Analytic Isomorphism: The mapping $\Phi$ was shown to preserve all essential operations – addition, multiplication, function composition, and limit transitions – thereby maintaining the complete mathematical structure while transitioning between nonlinear and classical frameworks.
-  3. 3.
-    Martingale Classification: Our framework provided a clear distinction between symmetric $G$-martingales (fair in all volatility scenarios) and asymmetric $G$-martingales (favorable but fair only in worst-case scenarios), revealing the fundamental structure of stochastic processes under volatility uncertainty.
-The parameterized control perspective developed herein offers more than theoretical elegance; it provides a practical computational methodology for working with $G$-expectations by reducing complex nonlinear problems to families of more manageable classical problems.
-This approach maintains full mathematical rigor while offering significant advantages for both theoretical analysis and numerical implementation.
-Looking forward, this isomorphism establishes a solid foundation for further research in nonlinear stochastic calculus, with particular promise for applications in mathematical finance where volatility uncertainty plays a crucial role in pricing and risk management.
-The framework naturally extends to more complex stochastic processes and offers a new paradigm for understanding model uncertainty in quantitative economics.
+This work has developed a parametric representation framework for $G$-expectations that provides a systematic methodology for stochastic analysis under model uncertainty.
+We have demonstrated that this approach enables a structured decomposition of complex nonlinear problems, maintaining mathematical rigor while enhancing practical applicability.
+The principal advantages of our approach can be summarized as follows:
+  • Analytical Efficiency: The clear separation of the computational process into distinct phases allows researchers to leverage the full power of classical stochastic analysis.
+This structure focuses the specific challenges of nonlinear expectations primarily on the establishment of uniform estimates, thereby streamlining the overall analysis.
+  • Enhanced Precision: By performing a unified norm estimation over the entire parameter space at the final stage, our method captures the interdependencies among parameters.
+This global approach has been shown to yield sharper results compared to methods relying on the composition of piecewise sublinear estimates.
+  • Conceptual Clarity: The distinction between linear computational operations and the final robust synthesis offers an intuitive understanding of complex nonlinear phenomena through the lens of parameterized classical processes, making the $G$-framework more accessible.
+  • Financial Applications: In mathematical finance, our framework provides a natural foundation for robust pricing and hedging under volatility uncertainty.
+The parametric representation aligns with stochastic control approaches used in portfolio optimization and derivative pricing, where different volatility scenarios correspond to alternative market models.
+Our parametric methodology offers remarkable analytical convenience for studying solution properties.
+Once well-posedness is established, various characteristics of solutions can be directly investigated through the parameterized family, with complexity concentrated only in the final supremum operation.
+This not only significantly reduces analytical complexity but also improves estimation precision.
+Moreover, the framework exhibits strong extensibility to other types of model uncertainty beyond volatility ambiguity.
+For instance, in the context of $G$-Lévy processes, our parametric approach can be similarly applied by constructing appropriate parameterized families of classical Lévy processes.
+This demonstrates the versatility of our methodology in handling diverse forms of Knightian uncertainty within a unified analytical framework.
+Building upon the established foundation of $G$-expectation theory, our parametric framework provides a complementary perspective that broadens its practical utility.
+The results demonstrate that a careful structuring of parameter dependence can yield both theoretical insights and computational benefits.
+Looking forward, this parametric approach opens several promising directions for future research.
+The framework establishes a bridge to computational implementation by reformulating nonlinear problems into tractable families of classical processes.
+This creates significant potential for developing efficient numerical schemes to compute $G$-expectations and solve related stochastic equations.
+Furthermore, connections with other advanced stochastic theories and applications in domains such as financial mathematics and engineering systems represent fertile ground for further investigation.
+In summary, this work contributes to the development of $G$-expectation theory by providing a structured framework that effectively combines the strengths of classical stochastic analysis with the robustness requirements of model uncertainty.
+It is our hope that this perspective will facilitate both theoretical advances and practical applications in this evolving field.